### v1

# Measuring quantum relative entropy with finite-size effect

Abstract
We study the estimation of relative entropy $D(\rho\|\sigma)$ when $\sigma$ is known. We show that the Cramér-Rao type bound equals the relative varentropy. Our estimator attains the Cramér-Rao type bound when the dimension $d$ is fixed. It also achieves the sample complexity $O(d^{2})$ when the dimension $d$ increases. This sample complexity is optimal when $\sigma$ is the complexity mixed state. Also, it has time complexity $O(d^{5}\mathop{\rm polylog}\nolimits d)$. Our proposed estimator unifiedly works under both settings.
I Introduction
Distinguishing a given known state $\sigma$ from another state $\rho$ is a fundamental task in quantum information. When the state $\rho$ is known as well as $\sigma$ and $n$ copies of the unknown state are prepared, quantum Stein’s lemma [1, 2] guarantees that the quantum relative entropy $D(\rho\|\sigma):=\mathrm{Tr}\rho(\log\rho-\log\sigma)$ gives the limitation of the optimal discrimination power. The measurement achieving the optimal performance depends only on the state $\sigma$ [3]. Further, quantum Sanov theorem [5, 4] shows that the construction of the test achieving the optimal performance depends only on the state $\sigma$ and the value of the quantum relative entropy $D(\rho\|\sigma)$. That is, when we construct the test achieving the optimal performance, the information of the quantum relative entropy $D(\rho\|\sigma)$ is needed in addition to the information of $\sigma$. Although the papers [6, 7] studied the estimation of von Neumann entropy and Rényi entropy, they did not study the estimation of the quantum relative entropy $D(\rho\|\sigma)$ with the known state $\sigma$. Due to the non-commutativity between $\rho$ and $\sigma$, the method in [6, 7] cannot be applied to the estimation of the quantum relative entropy $D(\rho\|\sigma)$ even when the state $\sigma$ is known. In particular, the simple use of Schur sampling [8, 9, 10, 11, 12, 13] does not work for the estimation of the quantum relative entropy $D(\rho\|\sigma)$ although the paper [6] employs Schur sampling. Therefore, this paper studies how to estimate the quantum relative entropy $D(\rho\|\sigma)$ with $n$ copies of the unknown state $\rho$ when $\sigma$ is known.
The estimation of $D(\rho\|\sigma)$ with the unknown state $\rho$ is formulated as a special case of the estimation of a function output with nuisance parameter, which was formulated in [14, Section 4.3]. In this formulation, the optimal performance of the estimation precision is given by the Cramér-Rao type bound with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case. As the first step, using the results for estimation with nuisance parameters, we show that the Cramér-Rao type bound of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing. The meaning of the Cramér-Rao type bound is the following; when $n$ copies of the unknown state $\rho$ is prepared, the mean square of our best estimator behaves as $\frac{V(\rho\|\sigma)}{n}+o(\frac{1}{n})$. Under special models, the recent paper [15] showed the existence of an estimator whose $n$ times the mean square error equals the Cramér-Rao type bound without nuisance parameter. However, such an estimator does not exist in general.
In practice, it is necessary to find an estimator to achieve the above performance. In addition, it is also desired to clearly evaluate the difference between its mean square error and the Cramér-Rao type bound. Since our problem is one-parameter estimation, it is possible to achieve the above bound without the use of collective measurements across several quantum systems. However, this method requires an adaptive choice of measurement. One typical method for this approach is a two-step estimator, which is composed of the pre-estimation stage and the local optimal estimation. When we have $n$ copies of the unknown state $\rho$, theoretically, in order to achieve the optimal bound, it is sufficient to allocate $\sqrt{n}$ copies to the pre-estimation stage [17, 16]. Although this method achieves the optimal mean square error in the order $\frac{1}{n}$, i.e., the first order, it is not so clear how the mean square error is close to the leading term $\frac{V(\rho\|\sigma)}{n}$. From the practical viewpoint, it is needed to upper bound this difference by using the number $n$ of unknown copies. Unfortunately, the above method has difficulty to evaluate the difference between its mean square error and $\frac{V(\rho\|\sigma)}{n}$. The papers [18, 19] also proposed different adaptive methods for the parameter estimation, but their methods have the same problem. In addition, the above adaptive approach does not work when the dimension $d$ of the quantum system increases in addition to the number $n$ of copies.
In summary, considering the finite size effect of our estimation, we need to find an estimator of the quantum relative entropy $D(\rho\|\sigma)$ to satisfy the following conditions. It asymptotically attains the Cramér-Rao type bound $V(\rho\|\sigma)$ in the sense of the order $\frac{1}{n}$ when the dimension $d$ is fixed. The difference between its mean square error and the Cramér-Rao type bound is evaluated with the number $n$ of copies. Further, the estimator works well even when the dimension $d$ of the quantum system increases in addition to the number $n$ of the prepared copies.
As the second step, to satisfy the above requirements, we propose an alternative estimation method by using Schur transform [20]. Then, we evaluate its mean square error and clarify its difference from the Cramér-Rao type bound with a finite number $n$. In addition, we evaluate the tail probability.
As the third step, similar to the papers [6, 7], we study the case when the dimension $d$ increases when $n$ increases. Unlike the estimation of von Neumann entropy, the range of the quantum relative entropy nor the relative varentropy is not bounded even with fixed dimension $d$. Therefore, to study the sample complexity, we need to restrict the possible range of $\rho$ and $\sigma$. We show that the sample complexity of our method is $O(d^{2})$ under a certain natural restriction for this range. In particular, the restriction of the possible range of $\sigma$ is crucial in this analysis. We also show that our time complexity is $O(d^{5}\mathop{\rm polylog}\nolimits d)$.
When $\sigma$ is fixed to the completely mixed state, our problem is the same as the estimation of von Neumann entropy, which was studied in the papers [6, 7]. Our result contains the case studied in the papers [6, 7]. In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ when our measurement is limited to Schur sampling. The paper [7] also showed that the lower bound of the sample complexity is $O(d)$ without this restriction. In this special case, we show that our measurement can be restricted to Schur sampling so that the lower bound by [6] holds even without this restriction.
The remaining part of this paper is organized as follows. First, Section II derives the Cramér-Rao type bound for the estimation of quantum relative entropy. Section III presents our estimator for quantum relative entropy, and its performance. Then, this section shows that this estimator asymptotically attains the Cramér-Rao type bound when $d$ is fixed. Section IV studies the sample complexity of our method when $d$ increases. Section V studies the time complexity of our method when $d$ increases. Section VI shows the optimality of the use of Schur sampling when $\sigma$ is fixed to the completely mixed state. Appendices are devoted to the proofs of the results presented in Section III.
II Cramér-Rao type bound
To clarify the minimum mean square error with the first order $\frac{1}{n}$, we calculate the Cramér-Rao type bound with nuisance parameters of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$. For this aim, we review the Cramér-Rao type bound for a general function $f(\rho)$ in the full parameter model under a $d$-dimensional system as follows. We choose a parametrization $\rho_{\theta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{d^{2}-1})\in\mathbb{R}^{d^{2}-1}$ such that the first parameter gives the value of the function to be estimated and the other parameters are orthogonal to the first parameter at $\rho=\rho_{*}$. That is, our parameter needs to satisfy the following conditions.
$\rho_{\theta_{0}}$ $=\rho_{*},$ (1)
$\frac{\partial f(\rho_{\theta})}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$ $=1,$ (2)
$\frac{\partial f(\rho_{\theta})}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}$ $=0$ (3)
for $j=2,\ldots,d^{2}-1$, where $\theta_{0}:=(\theta^{1}_{0},0,\ldots,0)$, and $\theta^{1}_{0}:=f(\rho_{*})$.
For the latter discussion, we introduce the symmetric logarithmic derivative (SLD) inner product. Given two Hermitian matrices $X,Y$, we define $X\circ Y:=(XY+YX)/2$. Then, we define the SLD inner product $\langle X,Y\rangle_{\rho}:=\mathrm{Tr}(\rho\circ X)Y=\mathrm{Tr}\rho(X\circ Y)$. We define the SLD norm $\|X\|_{\rho}:=\sqrt{\langle X,X\rangle_{\rho}}$. We choose the SLD $L_{1},\ldots,L_{d^{2}-1}$ as
$\frac{\partial\rho_{\theta}}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}=\rho_{*}\circ L_{j}$ (4)
for $j=1,\ldots,d^{2}-1$. We also assume the following orthogonal condition;
$\langle L_{1},L_{j}\rangle_{\rho_{*}}=\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}L_{1}=0$ (5)
for $j=2,\ldots,d^{2}-1$. Then, we can calculate the Cramér-Rao type bound, i.e., the minimum mean square error under the local unbiasedness condition even under the full parameter model.
Proposition 1 ([14, Theorem 5.3]).
Assume the conditions (1)–(5). The Cramér-Rao type bound for the estimation of $f(\rho_{\theta})$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}$ at $\theta=\theta_{0}$.
To calculate the Cramér-Rao type bound for $D(\rho\|\sigma)$ at $\rho=\rho_{*}$, we choose the Hermitian matrix
$L^{1}:=((\log\rho_{*}-\log\sigma)-D(\rho_{*}\|\sigma)I),$ (6)
and $d^{2}-1$ linearly independent Hermitian traceless matrices $X_{1},X_{2},\ldots X_{d^{2}-1}$ such that
$X_{1}$ $=V(\rho_{*}\|\sigma)^{-1}\rho_{*}\circ L^{1}$ (7)
$\mathrm{Tr}X_{j}L^{1}$ $=0$ (8)
for $j=2,\ldots,d^{2}-1$. Then, we choose our parametrization $\rho_{\theta}$ as
$\rho_{\theta}:=\rho_{*}+(\theta^{1}-\theta^{1}_{0})X_{1}+\sum_{j=2}^{d^{2}-1}\theta^{j}X_{j}.$ (9)
This parameterization satisfies the condition (1). Since $\frac{\partial\rho_{\theta}}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}=X_{j}$, we have
$\frac{\partial D(\rho_{\theta}\|\sigma)}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
$=$ $\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}\rho_{*}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\mathrm{Tr}X_{1}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}X_{1}$
=$\stackrel{{\scriptstyle(b)}}{{=}}$ $\mathrm{Tr}X_{1}L^{1}+\mathrm{Tr}X_{1}D(\rho_{*}\|\sigma)I+\mathrm{Tr}X_{1}\stackrel{{\scriptstyle(c)}}{{=}}1.$ (10)
Here, $(b)$ and $(c)$ follows from (6) and (7) and the relation $\langle L^{1},L^{1}\rangle_{\rho_{*}}=V(\rho_{*}\|\sigma)$, respectively. In addition, $(a)$ can be shown as follows.
$\mathrm{Tr}\rho_{\theta_{0}}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}\stackrel{{\scriptstyle(a)}}{{=}}\mathrm{Tr}\int_{0}^{1}\rho_{\theta_{0}}^{s}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}\rho_{\theta_{0}}^{1-s}ds$
$=$ $\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}=\mathrm{Tr}X_{1},$ (11)
where $(a)$ follows from [21, Eq. (6.32)].
Using (8) and a similar discussion, we have
$\frac{\partial D(\rho_{\theta}\|\sigma)}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}=\mathrm{Tr}X_{j}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}X_{j}=0.$ (12)
Then, (11) and (12) imply the conditions (2) and (3).
Due to (7) and (8), we have $L_{1}=V(\rho_{*}\|\sigma)^{-1}L^{1}$ and
$\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{j}}L_{1}=\mathrm{Tr}X_{j}L_{1}=0\hbox{ for }j=2,\ldots,d^{2}-1,$ (13)
which implies the condition (5). Therefore, since all the conditions (1) – (5) hold, due to Proposition 1, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}=V(\rho_{*}\|\sigma)^{-1}$ at $\theta=\theta_{0}$. That is, the minimum mean square error under the local unbiasedness condition is $V(\rho_{*}\|\sigma)$. That is, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is the relative varentropy $V(\rho_{*}\|\sigma)$.
III Estimator based on Schur transform
Similar to the papers [22, 3, 8, 9, 10, 11, 12, 13, 4, 23, 24], we employ Schur duality of ${\cal H}^{\otimes n}$. For this aim, we prepare several notations. A sequence of monotone-increasing non-negative integers $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{d})$ is called Young index. Although many references [24, 26, 25] define Young index as monotone-decreasing non-negative integers, we define it in the opposite way for notational convenience. We denote the set of Young indices $\bm{\lambda}$ with condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [24, Section 6.2], we have
${\cal H}^{\otimes n}=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}\otimes{\cal V}_{\bm{\lambda}}.$ (14)
Here, ${\cal U}_{\bm{\lambda}}$ expresses the irreducible space of $\mathop{\rm SU}(d)$ and ${\cal V}_{\bm{\lambda}}$ expresses the irreducible space of the representation $\pi$ of the permutation group $\mathfrak{S}_{n}$. We define $d_{\bm{\lambda}}:=\dim{\cal U}_{\bm{\lambda}}$, $d_{\bm{\lambda}}^{\prime}:=\dim{\cal V}_{\bm{\lambda}}$. Then, for states $\sigma$, $\rho$, we have
$\rho^{\otimes n}$ $=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (15)
$\sigma^{\otimes n}$ $=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (16)
where $\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits}$ is the completely mixed state on ${\cal V}_{\bm{\lambda}}$. We choose a basis $\{|u_{\bm{\lambda},j}\rangle\}_{j=1}^{d_{\bm{\lambda}}}$ of ${\cal U}_{\bm{\lambda}}$ such that $|u_{\bm{\lambda},j}\rangle$ is an eigenvector of $\sigma_{\bm{\lambda}}$.
To estimate the quantum relative entropy $D(\rho\|\sigma)$, we employ the POVM $M=\{M_{\bm{\lambda},j}\}_{\bm{\lambda},j}$ defined as
$M_{\bm{\lambda},j}:=|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes I({\cal V}_{\bm{\lambda}}),$ (17)
where $I({\cal V}_{\bm{\lambda}})$ is the identity operator on ${\cal V}_{\bm{\lambda}}$. The measurement outcome is given as $(\bm{\lambda},j)$ so that it obeys the probability distribution
$P_{\rho,\sigma}^{(n)}(\bm{\lambda},j):=\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}.$ (18)
To estimate the quantum relative entropy $D(\rho\|\sigma)$, we employ the estimate $x_{\bm{\lambda},j}$ defined as $-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$. This measurement can be done by Schur transform efficiently [20].
To evaluate the mean square error of this estimator, we denote the dimension of $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$ by $d_{n}$. Since as shown in [24, (6.16) and (6.18)], we have
$\dim{\cal U}_{\bm{\lambda}}\leq(n+1)^{d(d-1)/2},\quad|Y_{d}^{n}|\leq(n+1)^{d-1},$ (19)
the relation
$d_{n}\leq(n+1)^{(d+2)(d-1)/2}$ (20)
holds. The mean square error of this estimator is evaluated as follows.
Theorem 2.
For any states $\rho$ and $\sigma$, we have
$\mathop{\rm MSE}\nolimits_{n}(\rho\|\sigma):=\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)(x_{\bm{\lambda},j}-D(\rho\|\sigma))^{2}$
$\leq$ $\Big{(}\frac{1}{\sqrt{n}}\sqrt{V(\rho\|\sigma)}+\frac{1}{n}\log d_{n}\Big{)}^{2}.$ (21)
This theorem is proved in Appendix A. The difference between the upper bound (21) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}\log d_{n}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n}$. Using (20), this different is upper bounded by $\frac{1}{n^{2}}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\frac{(d+2)(d-1)}{2}\log(n+1)$. $n$ times of these terms go to zero.
Next, we focus on the tail probability for our estimator, i.e., we study the estimation error probability as
$\delta_{\rho,\sigma}^{(n),+,\epsilon}$ $:=P_{\rho,\sigma}^{(n)}\{(\bm{\lambda},j):x_{\bm{\lambda},j}-D(\rho\|\sigma)>\epsilon\}$ (22)
$\delta_{\rho,\sigma}^{(n),-,\epsilon}$ $:=P_{\rho,\sigma}^{(n)}\{(\bm{\lambda},j):x_{\bm{\lambda},j}-D(\rho\|\sigma)<-\epsilon\}.$ (23)
When $\epsilon$ behaves as $c/\sqrt{n}$, as shown in Appendix B the probabilities $\delta_{\rho,\sigma}^{(n),\pm,c/\sqrt{n}}$ converges to a normal distribution as follows.
Theorem 3.
For any states $\rho$ and $\sigma$, we have
[…]
This property is called local asymptotic normality.
When the error term $\epsilon$ is constant, the tail probability goes to zero exponentially. We use the sandwich relative entropy $D_{\alpha}(\sigma\|\rho):=\frac{1}{\alpha-1}\log\mathrm{Tr}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha}$. As shown in Appendix C, using the method in Section X of [23], we have the following theorem.
Theorem 4.
For any states $\rho$, $\sigma$, and $R,r>0$, we have
$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $d_{n}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (26)
$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n}e^{-nr}.$ (27)
In (26), $\alpha\in(0,1)$, and in (27), $\alpha>0$.
Theorem 4 guarantees that $\delta_{\rho,\sigma}^{(n),+,\epsilon}$ and $\delta_{\rho,\sigma}^{(n),-,\epsilon}$ go to zero exponentially. This property is called the large deviation.
IV Analysis with increasing $d$
Similar to [6, 7], we consider the case when the dimension $d$ increases. When $n=O(d^{4+\delta})$, $\frac{1}{\sqrt{n}}\log d_{n}\leq\frac{1}{\sqrt{n}}\frac{(d+1)(d-1)}{2}\log(n+1)\to 0$. In this case, due to Theorem 2, the Cramér-Rao type bound can be attained because the second term of RHS in (21) goes to zero.
Now, we consider when the error goes to zero under the condition $d\to\infty$. We denote the $d$-dimensional system by ${\cal H}_{d}$. To discuss this case, we consider the sequence of states $\sigma_{d}$ depending on the dimension $d$. We also choose a subset ${\cal S}_{d}$ of the set ${\cal S}({\cal H}_{d})$ of states on ${\cal H}_{d}$. We assume the following condition for $\sigma_{d}$ and ${\cal S}_{d}$;
$c_{0}:=\lim_{d\to\infty}\frac{1}{d^{2}}\max_{\rho\in{\cal S}_{d}}V(\rho\|\sigma_{d})<\infty.$ (28)
For example, the condition (28) holds under certain conditions.
Lemma 5.
When the minimum eigenvalue of $\sigma_{d}$ is lower bounded by $e^{-td}$ with a fixed real number $t>0$, the condition (28) holds even when ${\cal S}_{d}$ equals ${\cal S}({\cal H}_{d})$.
Proof.
Considering the definition of $V(\rho\|\sigma_{d})$ and the norm $\|~{}\|_{\rho}$, we have
$\sqrt{V(\rho\|\sigma_{d})}$
$=$ $\|\log\rho-(\mathrm{Tr}\rho\log\rho)-\log\sigma_{d}+(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}$
$\leq$ $\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}+\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}.$ (29)
Using the property of the norm $\|~{}\|_{\rho}$, we have
$\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}^{2}=\mathrm{Tr}\rho(\log\rho-(\mathrm{Tr}\rho\log\rho))^{2}$
$=$ $(\mathrm{Tr}\rho(\log\rho)^{2})-(\mathrm{Tr}\rho\log\rho)^{2}\leq(\mathrm{Tr}\rho(\log\rho)^{2})\stackrel{{\scriptstyle(a)}}{{\leq}}(\log d)^{2},$ (30)
where $(a)$ follows from [3, Lemma 8]. Similarly, we have
$\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}^{2}=\mathrm{Tr}\rho(\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d}))^{2}$
$=$ $(\mathrm{Tr}\rho(\log\sigma_{d})^{2})-(\mathrm{Tr}\rho\log\sigma_{d})^{2}\leq(\mathrm{Tr}\rho(\log\sigma_{d})^{2})\leq(td)^{2}.$ (31)
Combining (29), (30), and (31), we obtain the condition (28). ∎
Theorem 6.
When the condition (28) holds and $n=cd^{2}$, we have
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon}+\delta_{\rho,\sigma}^{(n),-,\epsilon}\leq$ $\frac{1}{\epsilon^{2}}\Big{(}\frac{\sqrt{c_{0}}}{\sqrt{c}}+\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}\Big{)}^{2}$
$\leq$ $\frac{(\sqrt{c_{0}}+4)^{2}}{c\epsilon^{2}}.$ (32)
When we choose $c$ to be $\frac{(\sqrt{c_{0}}+4)^{2}}{\epsilon^{\prime}\epsilon^{2}}$ for any $\epsilon,\epsilon^{\prime}>0$, we have
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon}+\delta_{\rho,\sigma}^{(n),-,\epsilon}\leq\epsilon^{\prime}.$ (33)
Hence, we obtain an upper bound $O(d^{2})$ of the sample complexity for the estimation of $D(\rho\|\sigma_{d})$ when $\rho$ belongs to ${\cal S}_{d}$.
Proof.
We choose $s\in(0,1)$. Using Weyl’s dimension formula [25, Theorem 7.1.7], we have
$\log\dim{\cal U}_{\bm{\lambda}}=\log\prod_{1\leq i<j\leq d}\frac{j-i+\lambda_{j}-\lambda_{i}}{j-i}$
$=$ $\!\!\sum_{1\leq i<j\leq d}\!\!\log\Big{(}1+\frac{\lambda_{j}-\lambda_{i}}{j-i}\Big{)}=\sum_{l=1}^{d}\sum_{i=1}^{d-l}\log\Big{(}1+\frac{\lambda_{l+i}-\lambda_{i}}{l}\Big{)}$
$=$ $\sum_{l=1}^{d}(d-l)\sum_{i=1}^{d-l}\frac{1}{d-l}\log\Big{(}1+\frac{\lambda_{l+i}-\lambda_{i}}{l}\Big{)}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\sum_{i=1}^{d-l}\frac{1}{d-l}\frac{\lambda_{l+i}-\lambda_{i}}{l}\Big{)}$
$\leq$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\frac{n}{(d-l)l}\Big{)}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\sum_{l=1}^{d}\frac{(d-l)}{s}\Big{(}\frac{n}{(d-l)l}\Big{)}^{s}=\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}},$ (34)
where $(a)$ follows from Jensen inequality for the concave function $x\mapsto\log(1+x)$ and the uniform distribution over $\{1,\ldots,d-l\}$. $(b)$ can be shown as follows.
$\log\Big{(}1+\frac{n}{(d-l)l}\Big{)}=\frac{1}{s}\log\Big{(}1+\frac{n}{(d-l)l}\Big{)}^{s}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\frac{1}{s}\log\Big{(}1+\Big{(}\frac{n}{(d-l)l}\Big{)}^{s}\Big{)}\stackrel{{\scriptstyle(d)}}{{\leq}}\frac{1}{s}\Big{(}\frac{n}{(d-l)l}\Big{)}^{s},$ (35)
where $(c)$ follows from the inequality $(x+y)^{s}\leq x^{s}+y^{s}$ for $x,y\geq 0$, and $(d)$ follows from the inequality $\log(1+x)\leq x$. Thus, we have
$\log d_{n}\leq\log|Y_{d}^{n}|\max_{\bm{\lambda}}\dim{\cal U}_{\bm{\lambda}}$
$\leq$ $\log|Y_{d}^{n}|+\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}}$
$\leq$ $(d-1)\log(n+1)+\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}}.$ (36)
When $n=cd^{2}$, we have
$\frac{1}{d^{2}}\log\dim{\cal U}_{\bm{\lambda}}\leq\frac{1}{d^{2}}\sum_{l=1}^{d}\frac{(d-l)^{1-s}c^{s}d^{2s}}{sl^{s}}$
[…]
Since the term $\frac{1}{n}(d-1)\log(n+1)$ goes to zero, the term $\frac{1}{n}\log d_{n}$ is evaluated as
$\lim_{n\to\infty}\frac{1}{n}\log d_{n}\leq\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}.$ (38)
The definition of $c_{0}$ implies
$\lim_{n\to\infty}\frac{1}{\sqrt{n}}\sqrt{V(\rho\|\sigma)}=\frac{\sqrt{c_{0}}}{\sqrt{c}}.$ (39)
Thus, we have
$\lim_{n\to\infty}\mathop{\rm MSE}\nolimits_{n}(\rho\|\sigma_{d})$
$\stackrel{{\scriptstyle(e)}}{{\leq}}$ $\Big{(}\frac{\sqrt{c_{0}}}{\sqrt{c}}+\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}\Big{)}^{2}\stackrel{{\scriptstyle(f)}}{{\leq}}\frac{(\sqrt{c_{0}}+4)^{2}}{c},$ (40)
where $(e)$ follows from the combination of (38), (39), and Theorem 2, $(f)$ follows by choosing $s=1/2$. Finally, applying Chebyshev’s inequality, we obtain the desired inequalities. ∎
The key point of this proof is the derivation (34). In particular, the derivation (35) is the heart in this part. The idea of this derivation comes from the proof of (12) given in Appendix IV of the paper [27], which gives the security analysis in information theoretic security.
V Analysis on time complexity
To implement our estimator, we need to apply suitable unitary gates before the measurement based on the computation basis. For simplicity, we consider the case when $\sigma$ is a diagonal matrix. For a general density $\sigma$, we need to apply a unitary diagonalizing the density $\sigma$ to each system ${\cal H}$ before the following procedure.
First, we apply the Schur transform. The references [20, 28, 29] proposed its quantum algorithm. We evaluate its prevision as follows. Given a required error $\epsilon>0$ and a target unitary $V$, we say that a unitary $U$ is an $\epsilon$-approximate of $V$ when any pure state $|\psi\rangle$ satisfies
$\|(U-V)|\psi\rangle\|\leq\epsilon.$ (41)
When the constructed unitary is an $\epsilon$-approximate of the Schur transform, the time complexity of the construction [20, 29] is $O(nd^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$. Since these constructions realize the approximation (41), the estimation error probability increases at most only $\epsilon^{2}$. We denote the distribution of $(\lambda,j)$ when we employ an $\epsilon$-approximate of the Schur transform by $P_{\rho,\sigma}^{(n),\epsilon}$.
In the classical part of our estimator, we need to calculate $x_{\bm{\lambda},j}=-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$. When $\sigma$ has diagonal elements $s_{1},\ldots,s_{d}$, the eigenvalue of $\sigma^{\otimes n}$ on the projection $M_{\bm{\lambda},j}$ has the form $\prod_{j=1}^{d}s_{j}^{n_{j}}$, where $\sum_{j=1}^{d}n_{j}=n$. Then, we have
$\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}=\dim{\cal V}_{\bm{\lambda}}\prod_{j=1}^{d}s_{j}^{n_{j}}.$ (42)
For a Young index $\bm{\lambda}$, the dimension ${\cal V}_{\bm{\lambda}}$ is evaluated as
$\dim{\cal V}_{\bm{\lambda}}$ $=\frac{n!}{\bm{\lambda}!}e(\bm{\lambda})$ (43)
$e(\bm{\lambda})$ $:=\prod_{j>i}\frac{\lambda_{j}-\lambda_{i}-i+j}{\lambda_{j}+j-i}<1,$ (44)
where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [26, Eq. (2.72)]. Using [21, Theorem 2.5], we have
$e^{nH(\frac{\bm{\lambda}}{n})}(n+1)^{-(d-1)}\leq\frac{n!}{\bm{\lambda}!}\leq e^{nH(\frac{\bm{\lambda}}{n})}.$ (45)
The value $e(\bm{\lambda})$ is evaluated as
$-\log e(\bm{\lambda})=\sum_{j>i}\log\frac{\lambda_{j}+j-i}{\lambda_{j}-\lambda_{i}-i+j}$
$=$ $\sum_{j>i}\log(1+\frac{\lambda_{i}}{\lambda_{j}-\lambda_{i}+j-i})\leq\sum_{j>i}\log(1+\frac{\lambda_{i}}{j-i})$
$=$ $\sum_{l=1}^{d}(d-l)\sum_{i=1}^{d-l}\frac{1}{d-l}\log\Big{(}1+\frac{\lambda_{i}}{l}\Big{)}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\sum_{i=1}^{d-l}\frac{1}{d-l}\frac{\lambda_{i}}{l}\Big{)}$
$\leq$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\frac{n}{(d-l)l}\Big{)}\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}},$ (46)
where $(a)$ follows from Jensen inequality for the concave function $x\mapsto\log(1+x)$, and $(b)$ follows from (35) in the same way as (34).
Instead of $x_{\bm{\lambda},j}$, we use its approximation $x_{\bm{\lambda},j}^{*}:=H(\frac{\bm{\lambda}}{n})-\sum_{j=1}^{d}\frac{n_{j}}{n}\log s_{j}$, whose time complexity is $O(n)$. Then, we have
$0\leq x_{\bm{\lambda},j}^{*}-x_{\bm{\lambda},j}\leq$ $\frac{1}{n}(d\log(n+1)-\log e(\bm{\lambda}))$
$\leq$ $\frac{1}{n}\Big{(}d\log(n+1)-\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}}\Big{)}.$ (47)
When $n=cd^{2}$, $\lim_{n\to\infty}\frac{1}{n}(d\log(n+1)-\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}})\leq\frac{4}{\sqrt{c}}$.
Given an arbitrary number $\epsilon>0$, we choose $c=\frac{16}{\epsilon}$. When $n=cd^{2}$, the limit of the upper bound in (47) can be bounded by $\epsilon$. Then, we define
$\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}$ $:=P_{\rho,\sigma}^{(n),\epsilon_{1}}\{(\bm{\lambda},j):x_{\bm{\lambda},j}^{*}-D(\rho\|\sigma)>\epsilon\}$ (48)
$\delta_{\rho,\sigma}^{(n),-,\epsilon,\epsilon_{1}}$ $:=P_{\rho,\sigma}^{(n),\epsilon_{1}}\{(\bm{\lambda},j):x_{\bm{\lambda},j}^{*}-D(\rho\|\sigma)<-\epsilon\}.$ (49)
Thus, when $n=cd^{2}$, using (33), we have
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}+\delta_{\rho,\sigma}^{(n),-,\epsilon,\epsilon_{1}}\leq\frac{(\sqrt{c_{0}}+4)^{2}}{c(\epsilon-\frac{4}{\sqrt{c}})^{2}}+\epsilon_{1}^{2}.$ (50)
Hence, when we choose $\epsilon_{1}=\sqrt{\epsilon^{\prime}/2}$ and $c=\min(\frac{64}{\epsilon^{2}},\frac{8(4+\sqrt{c_{0}})^{2}}{\epsilon^{2}\epsilon^{\prime}})$ for any $\epsilon,\epsilon^{\prime}>0$, we have
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}+\delta_{\rho,\sigma}^{(n),-,\epsilon,\epsilon_{1}}\leq\epsilon^{\prime}.$ (51)
Overall, when $n=O(d^{2})$, we use the construction of Schur transform by [20, 29] with the time complexity $O(d^{5}\mathop{\rm polylog}\nolimits(d))$ and the estimate $x_{\bm{\lambda},j}$ with the time complexity $O(n)=O(d^{2})$. That is, the construction has a total time complexity $O(d^{5}\mathop{\rm polylog}\nolimits(d))$.
VI Optimality of use of Schur sampling
The paper [6] studied the case when $\sigma_{d}$ is the completely mixed state $\rho_{\mathop{\rm mix}\nolimits}$. In this case, the paper [6] showed that the lower bound of sample complexity is $O(d^{2})$ when $d$ is increasing, Schur sampling is employed. In this case, the estimate $x_{\bm{\lambda},j}$ of our estimator depends only on $\bm{\lambda}$. So, $x_{\bm{\lambda},j}$ is given as a function of the measurement outcome of Schur sampling.
In the following, we show that the optimal estimator can be given with the use of Schur sampling when $\sigma_{d}$ is the completely mixed state $\rho_{\mathop{\rm mix}\nolimits}$. That is, for any estimator POVM $M$, there exists an estimator POVM $M^{\prime}$ based on Schur sampling such that the performance of $M^{\prime}$ is not worse than the performance of $M$. For its precise statement, we introduce the following minimax criterion. When the true state is $\rho$ and the estimator POVM is $M$, we denote the distribution of the estimate $\hat{x}$ of the value $D(\rho\|\rho_{\mathop{\rm mix}\nolimits})$ by $P_{\rho;M}$. We focus on the following quantities
$R(\rho,M,\epsilon):=\max_{g\in\mathop{\rm SU}(d)}P_{g\rho g^{\dagger};M}\{\hat{x}||x-D(\rho\|\rho_{\mathop{\rm mix}\nolimits})|>\epsilon\}.$ (52)
We denote the projection to ${\cal U}_{\bm{\lambda}}\otimes{\cal V}_{\bm{\lambda}}$ by $E_{\bm{\lambda}}$. Then, we have the following lemma.
Lemma 7.
For any estimator POVM $M=\{(M_{i},\hat{x}_{i})\}_{i}$, there exists an estimator POVM $M^{\prime}=\{(M^{\prime}_{i},\hat{x}_{i})\}_{i}$ such that each operator $M^{\prime}_{i}$ is written as the form
$M^{\prime}_{i}=\sum_{\bm{\lambda}}c_{i|\bm{\lambda}}E_{\bm{\lambda}},$ (53)
and
$R(\rho,M,\epsilon)\geq R(\rho,M^{\prime},\epsilon).$ (54)
Here, $\hat{x}_{i}$ expresses the estimate when a measurement outcome $i$ is observed.
Eq. (53) means that the estimator $M^{\prime}$ is given by using Schur sampling. Eq. (54) means that the estimator $M^{\prime}$ is better than the estimator $M$. Hence, if we do not care about the time complexity, it is allowed that our estimator is a function of Schur sampling including a randomized function of Schur sampling. Hence, combining the result by [6], we find that the lower bound of sample complexity is $O(d^{2})$ when $d$ is increasing and $\sigma$ is the completely mixed state.
Proof.
To prove this lemma, we employ the method of Holevo’s covariant measurement [30, Chapter 4], [24, Chapter 4]. We denote the invariant probability measure on $\mathop{\rm SU}(d)$ by $\mu$. Since $M^{\prime}_{i}:=\int_{\mathop{\rm SU}(d)}(g^{\otimes n})^{\dagger}M_{i}g^{\otimes n}\mu(dg)$ is invariant for the representation $g^{\otimes n}$, it has the form $\sum_{\bm{\lambda}}c_{i|\bm{\lambda}}E_{\bm{\lambda}}$. We set the set $B_{\epsilon}:=\{i||x-D(\rho\|\rho_{\mathop{\rm mix}\nolimits})|>\epsilon\}$. Then, we have
$R(\rho,M,\epsilon)=\max_{g\in\mathop{\rm SU}(d)}\sum_{i\in B_{\epsilon}}\mathrm{Tr}g^{\otimes n}\rho^{\otimes n}(g^{\otimes n})^{\dagger}M_{i}$
$\geq$ $\int_{\mathop{\rm SU}(d)}\sum_{i\in B_{\epsilon}}\mathrm{Tr}g^{\otimes n}\rho^{\otimes n}(g^{\otimes n})^{\dagger}M_{i}\mu(dg)$
$=$ $\sum_{i\in B_{\epsilon}}\mathrm{Tr}\rho^{\otimes n}\Big{(}\int_{\mathop{\rm SU}(d)}(g^{\otimes n})^{\dagger}M_{i}g^{\otimes n}\mu(dg)\Big{)}$
$=$ $\sum_{i\in B_{\epsilon}}\mathrm{Tr}\rho^{\otimes n}M^{\prime}_{i}=R(\rho,M^{\prime},\epsilon).$ (55)
VII Discussion
We have studied the estimation of quantum relative entropy $D(\rho\|\sigma)$ when the second argument $\sigma$ is known. Our estimator has a good performance in the following two senses. First, when the dimension $d$ is fixed, our estimator attains Cramér-Rao type bound asymptotically. Second, when the dimension $d$ increases, our estimator has sample complexity $O(d^{2})$ under natural assumptions. This sample complexity is optimal when $\sigma$ is the completely mixed state. Third, when $d$ is large, although the calculation of our estimator is difficult, we have an alternative estimator whose time complexity is $O(d^{5}\mathop{\rm polylog}\nolimits(d))$. The difference between this alternative estimator is our original estimator is zero asymptotically. In the above way, our estimator works well unifiedly, which is a unique feature and did not appear in existing studies [6, 7, 14].
Since our estimator assumes the knowledge of the form of the second argument $\sigma$, it is needed to develop an estimator that works without this knowledge. In this case, we need to prepare $n$ copies of the unknown state $\sigma$ as well. To attain the Cramér-Rao type bound with fixed $d$, it is sufficient to estimate the form of the density matrix $\sigma$. However, such a method has a large time complexity when $d$ increases. It is an interesting open problem to derive an efficient estimator of the quantum relative entropy in the above setting.
Acknowledgement
MH is supported in part by the National Natural Science Foundation of China (Grant No. 62171212). The author is grateful for Professor Kun Fang to a helpful discussion.
Appendix A Evaluation of mean square error: Proof of Theorem 2
We define the state $\rho_{n}:=\sum_{\bm{\lambda}}\rho_{\bm{\lambda}}$ on the system $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$. We choose an orthogonal basis $\{|v_{\bm{\lambda},j^{\prime}}\}_{j^{\prime}}$ of ${\cal V}_{\bm{\lambda}}$.
$\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)(x_{\bm{\lambda},j}-D(\rho\|\sigma))^{2}$
$=$ $\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)\Big{(}\frac{1}{n}(-\log\langle u_{\bm{\lambda},j}|\sigma_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle-D(\rho\|\sigma)\Big{)}^{2}$
$=$ $\sum_{(\bm{\lambda},j,j^{\prime})}\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
$\Big{(}\frac{1}{n}(-\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
$+\log\langle v_{\bm{\lambda},j^{\prime}}|\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits}|v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma)\Big{)}^{2}$
$=$ $\sum_{(\bm{\lambda},j,j^{\prime})}\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
$\Big{(}\frac{1}{n}(-\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
$+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sum_{\bm{\lambda}^{\prime}}I_{{\cal U}_{\bm{\lambda}}}\otimes\rho_{\bm{\lambda}^{\prime},\mathop{\rm mix}\nolimits}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)$
$-D(\rho\|\sigma)\Big{)}^{2}$
$=$ $\mathrm{Tr}\rho^{\otimes n}\Big{(}\frac{1}{n}(-\log\sigma^{\otimes n}+\sum_{\bm{\lambda}}\log(I_{{\cal U}_{\bm{\lambda}}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits}))$
$-D(\rho\|\sigma)\Big{)}^{2}$
$=$ $\mathrm{Tr}\rho^{\otimes n}\Big{(}\frac{1}{n}(-\log\sigma^{\otimes n}+\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits}))$
$-D(\rho\|\sigma)-\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{)}^{2}$
$=$ $\mathrm{Tr}\rho^{\otimes n}\Big{(}\frac{1}{n}(-\log\sigma^{\otimes n}+\log\rho^{\otimes n})-D(\rho\|\sigma)$
$-\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{)}^{2}$
$=$ $\Big{\|}\frac{1}{n}(-\log\sigma^{\otimes n}+\log\rho^{\otimes n})-D(\rho\|\sigma)$
$-\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{\|}_{\rho^{\otimes n}}^{2}$
$\leq$ $\Big{(}\Big{\|}\frac{1}{n}(-\log\sigma^{\otimes n}+\log\rho^{\otimes n})-D(\rho\|\sigma)\Big{\|}_{\rho^{\otimes n}}$
$+\Big{\|}\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{\|}_{\rho^{\otimes n}}\Big{)}^{2}.$ (56)
Also, we have
$\Big{\|}\frac{1}{n}(-\log\sigma^{\otimes n}+\log\rho^{\otimes n})-D(\rho\|\sigma)\Big{\|}_{\rho^{\otimes n}}$
$=$ $\mathrm{Tr}\rho^{\otimes n}\Big{(}\frac{1}{n}(-\log\sigma^{\otimes n}+\log\rho^{\otimes n})-D(\rho\|\sigma)\Big{)}^{2}$
$=$ $\frac{1}{n}V(\rho\|\sigma),$ (57)
and
$\Big{\|}-\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{\|}_{\rho^{\otimes n}}$
$=$ $\mathrm{Tr}\rho^{\otimes n}\Big{(}\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{)}^{2}$
$=$ $\frac{1}{n^{2}}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{n^{2}}(\log d_{n})^{2},$ (58)
where $(a)$ follows from [3, Lemma 8]. The combination of (56), (57), and (58) implies (21).
Appendix B Local asymptotic normality: Proof of Theorem 3
We use the methods [3, 32, 31].
$x_{\bm{\lambda},j}-D(\rho\|\sigma)=-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}-D(\rho\|\sigma)$
$=$ $-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log d_{\bm{\lambda}}^{\prime})-D(\rho\|\sigma)$
$=$ $-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
$-\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma)$
$-\frac{1}{n}(\log d_{\bm{\lambda}}^{\prime}+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)$
$=$ $\frac{1}{n}(-\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
$+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma)$
$-\frac{1}{n}\log\langle u_{\bm{\lambda},j}|\rho_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle.$ (59)
The random variable $\sqrt{n}(-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma))$ has additional variable $j^{\prime}$. The joint distribution of $(\bm{\lambda},j,j^{\prime})$ is given as $\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$. This random variable coincides with $\sqrt{n}(\frac{1}{n}(\log{\cal E}_{\sigma^{\otimes n}}(\rho^{\otimes n})-\log\sigma^{\otimes n})-D(\rho\|\sigma))$ under the state ${\cal E}_{\sigma^{\otimes n}}(\rho^{\otimes n})$ studied in Section V of [32]. The discussion in Section V of [32] showed that it converges to Gaussian distribution with variance $V(\rho\|\sigma)$ in probability as $n\to\infty$.
In addition, in the same way as (58), using [3, Lemma 8], we have
$\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)(\frac{1}{\sqrt{n}}\log\langle u_{\bm{\lambda},j}|\rho_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle)^{2}$
$=$ $\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)(\frac{1}{\sqrt{n}}\log P_{\rho,\sigma}^{(n)}(\bm{\lambda},j))^{2}$
$=$ $\frac{1}{n}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\leq\frac{1}{n}(\log d_{n})^{2}$
$\leq$ $\frac{1}{n}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}\to 0.$ (60)
Thus, the random variable $\frac{1}{\sqrt{n}}\log\langle u_{\bm{\lambda},j}|\rho_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle$ under the distribution $P_{\rho,\sigma}^{(n)}$ converges to zero in probability as $n\to\infty$. Therefore, the probabilities $\delta_{\rho,\sigma}^{(n),\pm,c/\sqrt{n}}$ can be asymptotically characterized by Gaussian distribution with variance $V(\rho\|\sigma)$.
Appendix C Large deviation evaluation
C.1 Preparation
We define the pinching map $\Gamma_{\rho,n}$ as
$\Gamma_{\sigma,n}(\rho^{\prime})$
::=:= $\sum_{\bm{\lambda}\in Y_{d}^{n},j}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))\rho^{\prime}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))$ (61)
for a general state $\rho^{\prime}$ on ${\cal H}^{\otimes n}$. Since the system ${\cal V}_{\bm{\lambda}}$ has no information, it is sufficient to handle the two states $\oplus_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}$ and $\oplus_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}$ over $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$. We define the map
$\kappa_{\sigma,n}(\rho^{\prime}):=$ $\sum_{\bm{\lambda}\in Y_{d}^{n},j}\mathrm{Tr}_{{\cal V}_{\bm{\lambda}}}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))\rho^{\prime}$
$\cdot(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))$ (62)
for a general state $\rho^{\prime}$ on ${\cal H}^{\otimes n}$. Due to the above-mentioned structure, the states $\Gamma_{\sigma,n}(\rho^{\otimes n})$ and $\sigma^{\otimes n}$ are commutative with each other.
We have the pinching inequality:
$\rho^{\prime}\leq d_{n}\Gamma_{\sigma,n}(\rho^{\prime}).$ (63)
In fact, applying the matrix monotone function to (63), for $\alpha\in(0,1)$, we have
$d_{n}^{\alpha-1}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $\Big{(}d_{n}\Gamma_{\sigma,n}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}^{\alpha-1}$
$\leq$ $\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}.$ (64)
Thus, we have
$\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{\alpha}(\sigma^{\otimes n})^{1-\alpha}$
$=$ $\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $\mathrm{Tr}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha}$
$=$ $\mathrm{Tr}\big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\big{)}$
$\cdot\big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}\big{)}$
$=$ $\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
$\cdot\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
$\cdot((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $d_{n}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $d_{n}^{1-\alpha}(\mathrm{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n},$ (65)
where Step $(a)$ follows from (64).
C.2 Proof of Theorem 4
For $\alpha\in(0,1)$, we have
$\sum_{(\bm{\lambda},j):x_{\bm{\lambda},j}<R}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}$
$=$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\sigma^{\otimes n})>e^{-nR}\}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\sigma^{\otimes n})>e^{-nR}\kappa_{\sigma,n}(\rho^{\otimes n})\}$
$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{\alpha}e^{n\alpha R}$
$=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}(\sigma^{\otimes n})^{\alpha}e^{n\alpha R}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
$=$ $d_{n}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)}.$ (66)
Here, $(b)$ follows from $\kappa_{\sigma,n}(\rho^{\otimes n})\leq I$, and $(c)$ follows from (65).
For $\alpha>0$, we have
$\sum_{(\bm{\lambda},j):x_{\bm{\lambda},j}>R}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}$
$=$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\sigma^{\otimes n})<e^{-nR}\}$
$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\sigma^{\otimes n})<e^{-nR}\}\{\kappa_{\sigma,n}(\rho^{\otimes n})>e^{-nr}\}$
$+\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\rho^{\otimes n})\leq e^{-nr}\}$
$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\sigma^{\otimes n})<e^{-nR+nr}\kappa_{\sigma,n}(\rho^{\otimes n})\}$
$+\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\rho^{\otimes n})\leq e^{-nr}\}$
$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
$=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\mathrm{Tr}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
$=$ $(\mathrm{Tr}(\sigma^{-\frac{\alpha}{2(1+\alpha)}}\rho\sigma^{-\frac{\alpha}{2(1+\alpha)}})^{1+\alpha}e^{-\alpha(R-r)})^{n}+d_{n}e^{-nr}$
$=$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n}e^{-nr}.$ (67)
$(d)$ follows from the monotonicity of sandwitch Rényi relative entropy.

### v2

@@
 It also achieves the sample complexity $O(d^{2})$ when the dimension $d$ increases.
-This sample complexity is optimal when $\sigma$ is the complexity mixed state.
-Also, it has time complexity $O(d^{5}\mathop{\rm polylog}\nolimits d)$.
+This sample complexity is optimal when $\sigma$ is the completely mixed state.
+Also, it has time complexity $O(d^{6}\mathop{\rm polylog}\nolimits d)$.
 Our proposed estimator unifiedly works under both settings.
+relative entropy; sample complexity; Cramer-Rao type bound; Schur composition
 I Introduction
@@
 Therefore, this paper studies how to estimate the quantum relative entropy $D(\rho\|\sigma)$ with $n$ copies of the unknown state $\rho$ when $\sigma$ is known.
-The estimation of $D(\rho\|\sigma)$ with the unknown state $\rho$ is formulated as a special case of the estimation of a function output with nuisance parameter, which was formulated in [14, Section 4.3].
-In this formulation, the optimal performance of the estimation precision is given by the Cramér-Rao type bound with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case.
-As the first step, using the results for estimation with nuisance parameters, we show that the Cramér-Rao type bound of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing.
-The meaning of the Cramér-Rao type bound is the following; when $n$ copies of the unknown state $\rho$ is prepared, the mean square of our best estimator behaves as $\frac{V(\rho\|\sigma)}{n}+o(\frac{1}{n})$.
-Under special models, the recent paper [15] showed the existence of an estimator whose $n$ times the mean square error equals the Cramér-Rao type bound without nuisance parameter.
+The estimation of $D(\rho\|\sigma)$ with the unknown state $\rho$ is formulated as a special case of the estimation of a function output with nuisance parameters.
+This general problem setting has been studied in various researches [14, 15, 16, 17, 18, 19].
+The reference [19, Section III] addressed the estimation problem of $D(\rho\|\sigma)$ as a special case, and showed that the estimation error is lower bounded by the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing [20, 21] while it did not show its attainability.
+Under a general setting, the reference [16, Section 4.3] showed that the optimal performance of the estimation precision is given by the Cramér-Rao type bound $C$ with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case.
+The meaning of the Cramér-Rao type bound $C$ is the following; when $n$ copies of the unknown state $\rho$ is prepared under a fixed model, the mean square of our best estimator behaves as $\frac{C}{n}+o(\frac{1}{n})$.
+When we have no nuisance parameter, the recent paper [22] showed the existence of an estimator whose mean square error (MSE) equals $\frac{C}{n}$ without additional term under special models.
 However, such an estimator does not exist in general.
-In practice, it is necessary to find an estimator to achieve the above performance.
-In addition, it is also desired to clearly evaluate the difference between its mean square error and the Cramér-Rao type bound.
+In this paper, as the first step, using the results for estimation with nuisance parameters in [16, Section 4.3], we show that the Cramér-Rao type bound $C$ of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which was obtained as a lower bound in [19].
+In practice, it is necessary to find an estimator to achieve the MSE $\frac{C}{n}+o(\frac{1}{n})$.
+In addition, it is also desired to clearly evaluate the difference between its mean square error and the Cramér-Rao type bound $\frac{C}{n}$.
 Since our problem is one-parameter estimation, it is possible to achieve the above bound without the use of collective measurements across several quantum systems.
@@
 When we have $n$ copies of the unknown state $\rho$, theoretically, in order to achieve the optimal bound, it is sufficient to allocate $\sqrt{n}$ copies to the pre-estimation stage [23, 15].
-Although this method achieves the optimal mean square error in the order $\frac{1}{n}$, i.e., the first order, it is not so clear how the mean square error is close to the leading term $\frac{V(\rho\|\sigma)}{n}$.
-From the practical viewpoint, it is needed to upper bound this difference by using the number $n$ of unknown copies.
-Unfortunately, the above method has difficulty to evaluate the difference between its mean square error and $\frac{V(\rho\|\sigma)}{n}$.
-The papers [18, 19] also proposed different adaptive methods for the parameter estimation, but their methods have the same problem.
+Although this method achieves the optimal mean square error in the order $\frac{1}{n}$, i.e., the first order, it is not so clear how the mean square error is close to the leading term $\frac{C}{n}$.
+From the practical viewpoint, it is necessary to upper bound this difference by using the number $n$ of unknown copies.
+Unfortunately, the above method has difficulty to evaluate the difference between its mean square error and $\frac{C}{n}$.
+The papers [24, 25] also proposed different adaptive methods for parameter estimation, but their methods have the same problem.
 In addition, the above adaptive approach does not work when the dimension $d$ of the quantum system increases in addition to the number $n$ of copies.
@@
 Further, the estimator works well even when the dimension $d$ of the quantum system increases in addition to the number $n$ of the prepared copies.
-As the second step, to satisfy the above requirements, we propose an alternative estimation method by using Schur transform [20].
+As the second step of this paper, to satisfy the above requirements, we propose an alternative estimation method by using Schur transform [26].
 Then, we evaluate its mean square error and clarify its difference from the Cramér-Rao type bound with a finite number $n$.
@@
 We show that the sample complexity of our method is $O(d^{2})$ under a certain natural restriction for this range.
+Our method extensively improves the sample complexity over the simple application of the full tomography [27].
 In particular, the restriction of the possible range of $\sigma$ is crucial in this analysis.
-We also show that our time complexity is $O(d^{5}\mathop{\rm polylog}\nolimits d)$.
+We also show that our time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits d)$.
 When $\sigma$ is fixed to the completely mixed state, our problem is the same as the estimation of von Neumann entropy, which was studied in the papers [6, 7].
 Our result contains the case studied in the papers [6, 7].
-In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ when our measurement is limited to Schur sampling.
+In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ for the Empirical Young Diagram (EYD) algorithm, which is given as the entropy of the diagonal state determined by the Young indecies.
+Although this case improves the sample complexity over the simple application of the full tomography, the improvement is incremental.
 The paper [7] also showed that the lower bound of the sample complexity is $O(d)$ without this restriction.
-In this special case, we show that our measurement can be restricted to Schur sampling so that the lower bound by [6] holds even without this restriction.
+In this special case, we show that our measurement can be restricted to Schur sampling.
 The remaining part of this paper is organized as follows.
@@
 Section V studies the time complexity of our method when $d$ increases.
-Section VI shows the optimality of the use of Schur sampling when $\sigma$ is fixed to the completely mixed state.
 Appendices are devoted to the proofs of the results presented in Section III.
 II Cramér-Rao type bound
-To clarify the minimum mean square error with the first order $\frac{1}{n}$, we calculate the Cramér-Rao type bound with nuisance parameters of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$.
-For this aim, we review the Cramér-Rao type bound for a general function $f(\rho)$ in the full parameter model under a $d$-dimensional system as follows.
+To clarify the minimum mean square error with the first order $\frac{1}{n}$, we discuss the Cramér-Rao type bound $C$ with nuisance parameters, which is mathematically formulated as a lower bound of MES under the local unbiasedness condition.
+The reference [19, (3.24), (3.52)] showed that the varentropy $V(\rho_{*}\|\sigma)$ is a lower bound under the estimation of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$.
+That is, they showed that
+$C\geq V(\rho_{*}\|\sigma),$ (1)
+and did not discuss whether the equality holds.
+In the following, we show the equality in (1).
+For this aim, we review the Cramér-Rao type bound $C$ for a general function $f(\rho)$ in the full parameter model under a $d$-dimensional system as follows.
 We choose a parametrization $\rho_{\theta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{d^{2}-1})\in\mathbb{R}^{d^{2}-1}$ such that the first parameter gives the value of the function to be estimated and the other parameters are orthogonal to the first parameter at $\rho=\rho_{*}$.
@@
 for $j=2,\ldots,d^{2}-1$. Then, we can calculate the Cramér-Rao type bound, i.e., the minimum mean square error under the local unbiasedness condition even under the full parameter model.
-Proposition 1 ([14, Theorem 5.3]).
+Proposition 1 ([16, Theorem 5.3]).
 Assume the conditions (2)–(6). The Cramér-Rao type bound for the estimation of $f(\rho_{\theta})$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}$ at $\theta=\theta_{0}$.
@@
 $=$ $\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}=\mathrm{Tr}X_{1},$ (12)
-where $(a)$ follows from [21, Eq. (6.32)].
+where $(a)$ follows from [28, Eq. (6.32)].
 Using (9) and a similar discussion, we have
@@
 That is, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is the relative varentropy $V(\rho_{*}\|\sigma)$.
+Thus, the quality in (1) holds.
 III Estimator based on Schur transform
@@
 Although many references [31, 33, 32] define Young index as monotone-decreasing non-negative integers, we define it in the opposite way for notational convenience.
-We denote the set of Young indices $\bm{\lambda}$ with condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [24, Section 6.2], we have
+We denote the set of Young indices $\bm{\lambda}$ with the condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [31, Section 6.2], we have
 ${\cal H}^{\otimes n}=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}\otimes{\cal V}_{\bm{\lambda}}.$ (15)
@@
 This measurement can be done by Schur transform efficiently [26].
-To evaluate the mean square error of this estimator, we denote the dimension of $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$ by $d_{n}$.
-Since as shown in [24, (6.16) and (6.18)], we have
+To evaluate the mean square error of this estimator, we denote the dimension of $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$ by $d_{n,d}$.
+Since as shown in [31, (6.16) and (6.18)], we have
 $\dim{\cal U}_{\bm{\lambda}}\leq(n+1)^{d(d-1)/2},\quad|Y_{d}^{n}|\leq(n+1)^{d-1},$ (20)
 the relation
-$d_{n}\leq(n+1)^{(d+2)(d-1)/2}$ (20)
+$d_{n,d}\leq(n+1)^{(d+2)(d-1)/2}$ (21)
 holds. The mean square error of this estimator is evaluated as follows.
@@
 $\mathop{\rm MSE}\nolimits_{n}(\rho\|\sigma):=\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)(x_{\bm{\lambda},j}-D(\rho\|\sigma))^{2}$
-$\leq$ $\Big{(}\frac{1}{\sqrt{n}}\sqrt{V(\rho\|\sigma)}+\frac{1}{n}\log d_{n}\Big{)}^{2}.$ (21)
+$\leq$ $\Big{(}\frac{1}{\sqrt{n}}\sqrt{V(\rho\|\sigma)}+\frac{1}{n}\log d_{n,d}\Big{)}^{2}.$ (22)
 This theorem is proved in Appendix A.
-The difference between the upper bound (21) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}\log d_{n}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n}$.
+The difference between the upper bound (22) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}\log d_{n,d}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n,d}$.
 Using (21), this different is upper bounded by $\frac{1}{n^{2}}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\frac{(d+2)(d-1)}{2}\log(n+1)$.
@@
 For any states $\rho$, $\sigma$, and $R,r>0$, we have
-$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $d_{n}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (26)
-$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n}e^{-nr}.$ (27)
+$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $d_{n,d}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (27)
+$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n,d}e^{-nr}.$ (28)
 In (27), $\alpha\in(0,1)$, and in (28), $\alpha>0$.
@@
 Similar to [6, 7], we consider the case when the dimension $d$ increases.
-When $n=O(d^{4+\delta})$, $\frac{1}{\sqrt{n}}\log d_{n}\leq\frac{1}{\sqrt{n}}\frac{(d+1)(d-1)}{2}\log(n+1)\to 0$.
+When $n=O(d^{4+\delta})$, $\frac{1}{\sqrt{n}}\log d_{n,d}\leq\frac{1}{\sqrt{n}}\frac{(d+1)(d-1)}{2}\log(n+1)\to 0$.
 In this case, due to Theorem 2, the Cramér-Rao type bound can be attained because the second term of RHS in (22) goes to zero.
@@
 $c_{0}:=\lim_{d\to\infty}\frac{1}{d^{2}}\max_{\rho\in{\cal S}_{d}}V(\rho\|\sigma_{d})<\infty.$ (29)
-For example, the condition (28) holds under certain conditions.
+For example, when $\sigma_{d}$ is fixed to the completely mixed state $\rho_{\mathop{\rm mix}\nolimits}$, $V(\rho\|\rho_{\mathop{\rm mix}\nolimits})$ equals the varentropy $\mathrm{Tr}\rho(\log\rho)^{2}-(\mathrm{Tr}\rho\log\rho)^{2}$.
+Since Lemma 8 of [3] showed that $\mathrm{Tr}\rho(\log\rho)^{2}\leq(\log d)^{2}$, the condition (29) holds even with ${\cal S}_{d}={\cal S}({\cal H}_{d})$ so that the term $\frac{1}{n}\log d_{n,d}$ is the major term in the RHS of (22).
+Further, the condition (29) holds with ${\cal S}_{d}={\cal S}({\cal H}_{d})$ under certain conditions for $\sigma_{d}$.
 Lemma 5.
@@
 Combining (30), (31), and (32), we obtain the condition (29). ∎
+The sample complexity is characterized as the following theorem.
 Theorem 6.
@@
 $\leq$ $\frac{(\sqrt{c_{0}}+4)^{2}}{c\epsilon^{2}}.$ (33)
-When we choose $c$ to be $\frac{(\sqrt{c_{0}}+4)^{2}}{\epsilon^{\prime}\epsilon^{2}}$ for any $\epsilon,\epsilon^{\prime}>0$, we have
+When we choose $c$ to be $\frac{(\sqrt{c_{0}}+4)^{2}}{\epsilon^{\prime}\epsilon^{2}}$ for any $\epsilon,\epsilon^{\prime}>0$, i.e., $n=\frac{(\sqrt{c_{0}}+4)^{2}}{\epsilon^{\prime}\epsilon^{2}}d^{2}$, we have
 $\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon}+\delta_{\rho,\sigma}^{(n),-,\epsilon}\leq\epsilon^{\prime}.$ (34)
-Hence, we obtain an upper bound $O(d^{2})$ of the sample complexity for the estimation of $D(\rho\|\sigma_{d})$ when $\rho$ belongs to ${\cal S}_{d}$.
-Proof.
-We choose $s\in(0,1)$. Using Weyl’s dimension formula [25, Theorem 7.1.7], we have
+Hence, when the condition (34) is imposed for the estimation error, we obtain an upper bound $O(d^{2})$ of the sample complexity for the estimation of $D(\rho\|\sigma_{d})$ when $\rho$ belongs to ${\cal S}_{d}$.
+Now, we compare the sample complexity of our method with the sample complexity of the simple application of full tomography [27, 34].
+In particular, we employ the error analysis on the trance norm [27] rather than other error criteria because the trance norm is most useful to evaluate the error of the relative entropy.
+Their method achieves the estimation error
+$\|\hat{\rho}-\rho\|_{1}\leq\epsilon$ (35)
+with high probability by sample complexity $O(d^{2}/\epsilon^{2})$, where $\hat{\rho}$ is the estimated density and $\rho$ is the true density.
+When the known density $\sigma_{d}$ is the completely mixed state, our problem is reduced to the estimation of von Neumann entropy $S(\rho)$.
+Fannes inequality implies the relation
+$|S(\hat{\rho})-S(\rho)|\leq\|\hat{\rho}-\rho\|_{1}\log d.$ (36)
+That is, to achieve the condition
+$|S(\hat{\rho})-S(\rho)|\leq\epsilon^{\prime},$ (37)
+we need the sample complexity $O(d^{2}(\log d)^{2}/{\epsilon^{\prime}}^{2})$.
+Although our method and the method by [6] improves the sample complexity over the above method, the degree of the improvement is incremental in this case.
+When the known state $\sigma_{d}$ is a general state to satisfy the condition of Lemma 5, i.e., the minimum eigenvalue of $\sigma_{d}$ is lower bounded by $e^{-td}$ with a fixed real number $t>0$, the part $\log\sigma_{d}$ has the operator norm $td$ at most.
+Hence, we have
+$|D(\hat{\rho}\|\sigma_{d})-D(\rho\|\sigma_{d})|$
+$\leq$ $|S(\hat{\rho})-S(\rho)|+|\mathrm{Tr}(\hat{\rho}-\rho)\log\sigma_{d}|$
+$\leq$ $\|\hat{\rho}-\rho\|_{1}\log d+\|\hat{\rho}-\rho\|_{1}\|\log\sigma_{d}\|$
+$\leq$ $\|\hat{\rho}-\rho\|_{1}(\log d+td).$ (38)
+That is, to achieve the condition
+$|D(\hat{\rho}\|\sigma_{d})-D(\rho\|\sigma_{d})|\leq\epsilon^{\prime},$ (39)
+we need the sample complexity $O(d^{2}(\log d+td)^{2}/{\epsilon^{\prime}}^{2})=O(d^{4}/{\epsilon^{\prime}}^{2})$.
+Our method greatly improves the sample complexity over the the simple application of full tomography [27].
+That is, our extension to the estimation of the relative entropy realizes significant improvement in the sample complexity over the above method.
+Proof of Theorem 6.
+We choose $s\in(0,1)$. Using Weyl’s dimension formula [32, Theorem 7.1.7], we have
 $\log\dim{\cal U}_{\bm{\lambda}}=\log\prod_{1\leq i<j\leq d}\frac{j-i+\lambda_{j}-\lambda_{i}}{j-i}$
@@
 where $(c)$ follows from the inequality $(x+y)^{s}\leq x^{s}+y^{s}$ for $x,y\geq 0$, and $(d)$ follows from the inequality $\log(1+x)\leq x$. Thus, we have
-$\log d_{n}\leq\log|Y_{d}^{n}|\max_{\bm{\lambda}}\dim{\cal U}_{\bm{\lambda}}$
+$\log d_{n,d}\leq\log|Y_{d}^{n}|\max_{\bm{\lambda}}\dim{\cal U}_{\bm{\lambda}}$
 $\leq$ $\log|Y_{d}^{n}|+\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}}$
@@
 […]
-Since the term $\frac{1}{n}(d-1)\log(n+1)$ goes to zero, the term $\frac{1}{n}\log d_{n}$ is evaluated as
-$\lim_{n\to\infty}\frac{1}{n}\log d_{n}\leq\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}.$ (38)
+Since the term $\frac{1}{n}(d-1)\log(n+1)$ goes to zero, the term $\frac{1}{n}\log d_{n,d}$ is evaluated as
+$\lim_{d\to\infty}\frac{1}{n}\log d_{n,d}\leq\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}.$ (44)
 The definition of $c_{0}$ implies
-$\lim_{n\to\infty}\frac{1}{\sqrt{n}}\sqrt{V(\rho\|\sigma)}=\frac{\sqrt{c_{0}}}{\sqrt{c}}.$ (39)
+$\lim_{d\to\infty}\frac{1}{\sqrt{n}}\max_{\rho\in{\cal S}_{d}}\sqrt{V(\rho\|\sigma_{d})}=\frac{\sqrt{c_{0}}}{\sqrt{c}}.$ (45)
 Thus, we have
-$\lim_{n\to\infty}\mathop{\rm MSE}\nolimits_{n}(\rho\|\sigma_{d})$
+$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\mathop{\rm MSE}\nolimits_{n}(\rho\|\sigma_{d})$
 $\stackrel{{\scriptstyle(e)}}{{\leq}}$ $\Big{(}\frac{\sqrt{c_{0}}}{\sqrt{c}}+\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}\Big{)}^{2}\stackrel{{\scriptstyle(f)}}{{\leq}}\frac{(\sqrt{c_{0}}+4)^{2}}{c},$ (46)
@@
 In particular, the derivation (41) is the heart in this part.
-The idea of this derivation comes from the proof of (12) given in Appendix IV of the paper [27], which gives the security analysis in information theoretic security.
+The idea of this derivation comes from the proof of (12) given in Appendix IV of the paper [35], which gives the security analysis in information-theoretic security.
 V Analysis on time complexity
@@
 For a general density $\sigma$, we need to apply a unitary diagonalizing the density $\sigma$ to each system ${\cal H}$ before the following procedure.
-First, we apply the Schur transform.
-The references [20, 28, 29] proposed its quantum algorithm.
+First, we apply the Schur transform $U_{Schur}$, which realizes a basis change to block-diagonalize $U^{\otimes n}$ for an arbitrary unitary $U$.
+The references [26, 36, 37, 38] proposed its quantum algorithm.
 We evaluate its prevision as follows.
@@
 $\|(U-V)|\psi\rangle\|\leq\epsilon.$ (47)
-When the constructed unitary is an $\epsilon$-approximate of the Schur transform, the time complexity of the construction [20, 29] is $O(nd^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$.
+When the constructed unitary is an $\epsilon$-approximate of the Schur transform $U_{Schur}$, the time complexity of the construction [37] is $O(nd^{4}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$.
 Since these constructions realize the approximation (47), the estimation error probability increases at most only $\epsilon^{2}$.
-We denote the distribution of $(\lambda,j)$ when we employ an $\epsilon$-approximate of the Schur transform by $P_{\rho,\sigma}^{(n),\epsilon}$.
+We denote the distribution of $(\lambda,j)$ when we employ an $\epsilon$-approximate of the Schur transform $U_{Schur}$ by $P_{\rho,\sigma}^{(n),\epsilon}$.
 In the classical part of our estimator, we need to calculate $x_{\bm{\lambda},j}=-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$.
@@
 $e(\bm{\lambda})$ $:=\prod_{j>i}\frac{\lambda_{j}-\lambda_{i}-i+j}{\lambda_{j}+j-i}<1,$ (50)
-where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [26, Eq. (2.72)]. Using [21, Theorem 2.5], we have
+where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [33, Eq. (2.72)]. Using [28, Theorem 2.5], we have
 $e^{nH(\frac{\bm{\lambda}}{n})}(n+1)^{-(d-1)}\leq\frac{n!}{\bm{\lambda}!}\leq e^{nH(\frac{\bm{\lambda}}{n})}.$ (51)
@@
 $\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}+\delta_{\rho,\sigma}^{(n),-,\epsilon,\epsilon_{1}}\leq\epsilon^{\prime}.$ (57)
-Overall, when $n=O(d^{2})$, we use the construction of Schur transform by [20, 29] with the time complexity $O(d^{5}\mathop{\rm polylog}\nolimits(d))$ and the estimate $x_{\bm{\lambda},j}$ with the time complexity $O(n)=O(d^{2})$.
-That is, the construction has a total time complexity $O(d^{5}\mathop{\rm polylog}\nolimits(d))$.
-VI Optimality of use of Schur sampling
-The paper [6] studied the case when $\sigma_{d}$ is the completely mixed state $\rho_{\mathop{\rm mix}\nolimits}$.
-In this case, the paper [6] showed that the lower bound of sample complexity is $O(d^{2})$ when $d$ is increasing, Schur sampling is employed.
-In this case, the estimate $x_{\bm{\lambda},j}$ of our estimator depends only on $\bm{\lambda}$.
-So, $x_{\bm{\lambda},j}$ is given as a function of the measurement outcome of Schur sampling.
-In the following, we show that the optimal estimator can be given with the use of Schur sampling when $\sigma_{d}$ is the completely mixed state $\rho_{\mathop{\rm mix}\nolimits}$.
-That is, for any estimator POVM $M$, there exists an estimator POVM $M^{\prime}$ based on Schur sampling such that the performance of $M^{\prime}$ is not worse than the performance of $M$.
-For its precise statement, we introduce the following minimax criterion.
-When the true state is $\rho$ and the estimator POVM is $M$, we denote the distribution of the estimate $\hat{x}$ of the value $D(\rho\|\rho_{\mathop{\rm mix}\nolimits})$ by $P_{\rho;M}$.
-We focus on the following quantities
-$R(\rho,M,\epsilon):=\max_{g\in\mathop{\rm SU}(d)}P_{g\rho g^{\dagger};M}\{\hat{x}||x-D(\rho\|\rho_{\mathop{\rm mix}\nolimits})|>\epsilon\}.$ (52)
-We denote the projection to ${\cal U}_{\bm{\lambda}}\otimes{\cal V}_{\bm{\lambda}}$ by $E_{\bm{\lambda}}$. Then, we have the following lemma.
-Lemma 7.
-For any estimator POVM $M=\{(M_{i},\hat{x}_{i})\}_{i}$, there exists an estimator POVM $M^{\prime}=\{(M^{\prime}_{i},\hat{x}_{i})\}_{i}$ such that each operator $M^{\prime}_{i}$ is written as the form
-$M^{\prime}_{i}=\sum_{\bm{\lambda}}c_{i|\bm{\lambda}}E_{\bm{\lambda}},$ (53)
-and
-$R(\rho,M,\epsilon)\geq R(\rho,M^{\prime},\epsilon).$ (54)
-Here, $\hat{x}_{i}$ expresses the estimate when a measurement outcome $i$ is observed.
-Eq.
-(53) means that the estimator $M^{\prime}$ is given by using Schur sampling.
-Eq.
-(54) means that the estimator $M^{\prime}$ is better than the estimator $M$.
-Hence, if we do not care about the time complexity, it is allowed that our estimator is a function of Schur sampling including a randomized function of Schur sampling.
-Hence, combining the result by [6], we find that the lower bound of sample complexity is $O(d^{2})$ when $d$ is increasing and $\sigma$ is the completely mixed state.
-Proof.
-To prove this lemma, we employ the method of Holevo’s covariant measurement [30, Chapter 4], [24, Chapter 4].
-We denote the invariant probability measure on $\mathop{\rm SU}(d)$ by $\mu$.
-Since $M^{\prime}_{i}:=\int_{\mathop{\rm SU}(d)}(g^{\otimes n})^{\dagger}M_{i}g^{\otimes n}\mu(dg)$ is invariant for the representation $g^{\otimes n}$, it has the form $\sum_{\bm{\lambda}}c_{i|\bm{\lambda}}E_{\bm{\lambda}}$.
-We set the set $B_{\epsilon}:=\{i||x-D(\rho\|\rho_{\mathop{\rm mix}\nolimits})|>\epsilon\}$.
-Then, we have
-$R(\rho,M,\epsilon)=\max_{g\in\mathop{\rm SU}(d)}\sum_{i\in B_{\epsilon}}\mathrm{Tr}g^{\otimes n}\rho^{\otimes n}(g^{\otimes n})^{\dagger}M_{i}$
-$\geq$ $\int_{\mathop{\rm SU}(d)}\sum_{i\in B_{\epsilon}}\mathrm{Tr}g^{\otimes n}\rho^{\otimes n}(g^{\otimes n})^{\dagger}M_{i}\mu(dg)$
-$=$ $\sum_{i\in B_{\epsilon}}\mathrm{Tr}\rho^{\otimes n}\Big{(}\int_{\mathop{\rm SU}(d)}(g^{\otimes n})^{\dagger}M_{i}g^{\otimes n}\mu(dg)\Big{)}$
-$=$ $\sum_{i\in B_{\epsilon}}\mathrm{Tr}\rho^{\otimes n}M^{\prime}_{i}=R(\rho,M^{\prime},\epsilon).$ (55)
-VII Discussion
+Overall, when $n=O(d^{2})$, we use the construction of Schur transform $U_{Schur}$ by [37] with the time complexity $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$ and the estimate $x_{\bm{\lambda},j}$ with the time complexity $O(n)=O(d^{2})$.
+That is, the total construction has a total time complexity $O(d^{6}\mathop{\rm polylog}\nolimits(d))$ to realize the condition (57) with a fixed value $\epsilon>0$.
+Remark 7.
+Initially, the paper [26] gives a polynomial algorithm for Schur transform $U_{Schur}$.
+However, it did not give a precise evaluation of its time complexity.
+Although the paper [37] writes that the protocol by the paper [26] has time complexity $O\left(nd^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon)\right)$ in its Table 1, the paper [37] does not give the derivation of this evaluation because its main focus is the implementation of the mixed Schur transform $U_{MixedSchur}$, which realizes a basis change to block-diagonalize $U^{\otimes n}\otimes\overline{U}^{\otimes m}$ for an arbitrary unitary $U$.
+In fact, the paper [37] proved that its time complexity is $O\left((n+m)d^{4}\mathop{\rm polylog}\nolimits(n,m,d,1/\epsilon)\right)$, and the recent paper [38] also proposed the same algorithm as the algorithm by [37].
+In this paper, to make a safer statement, we employ the statement that the time complexity of the unitary $U_{MixedSchur}$ is $O\left((n+m)d^{4}\mathop{\rm polylog}\nolimits(n,m,d,1/\epsilon)\right)$.
+Its special case with $m=0$ corresponds to $U_{Schur}$ and covers our case with the time complexity $O\left(nd^{4}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon)\right)$.
+Remark 8.
+The paper [7] also mentions that there exists an algorithm for Schur transform $U_{Schur}$ whose time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$ when $n=O(d^{2})$ [39].
+First, they employ the algorithm for weak Schur sampling given in [40, Section 4.2.2].
+Although this method employs the quantum Fourier transform the permutation group $\mathfrak{S}_{n}$, they proposed to use the algorithm by [41] for this part.
+The obtained algorithm has time complexity $O(n^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$.
+When $n=d^{2}$, it equals $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$.
+This implementation of weak Schur sampling was mentioned earlier in [42, Section 1.1.2].
+VI Discussion
 We have studied the estimation of quantum relative entropy $D(\rho\|\sigma)$ when the second argument $\sigma$ is known.
@@
 First, when the dimension $d$ is fixed, our estimator attains Cramér-Rao type bound asymptotically.
-Second, when the dimension $d$ increases, our estimator has sample complexity $O(d^{2})$ under natural assumptions.
-This sample complexity is optimal when $\sigma$ is the completely mixed state.
-Third, when $d$ is large, although the calculation of our estimator is difficult, we have an alternative estimator whose time complexity is $O(d^{5}\mathop{\rm polylog}\nolimits(d))$.
-The difference between this alternative estimator is our original estimator is zero asymptotically.
+Second, when the dimension $d$ increases, our estimator has sample complexity $O(d^{2})$ under natural assumptions, which greatly improves the sample complexity by the simple application of full tomography [27].
+Third, when $d$ is large, although the calculation of our estimator is difficult, we have an alternative estimator whose time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits(d))$.
+The difference between this alternative estimator and our original estimator approaches zero asymptotically.
 In the above way, our estimator works well unifiedly, which is a unique feature and did not appear in existing studies [6, 7, 16].
@@
 Acknowledgement
-MH is supported in part by the National Natural Science Foundation of China (Grant No. 62171212). The author is grateful for Professor Kun Fang to a helpful discussion.
+The author is supported in part by the National Natural Science Foundation of China (Grant No.
+62171212).
+The author is grateful for Professor Kun Fang to a helpful discussion.
+Also, the author would like to thank Mr Dmitry Grinko and Dr.
+Qisheng Wang for helpful discussions on the time complexity of Schur transform.
+In addition, the author would like to thank Prof.
+Jun Suzuki and Dr.
+Francesco Albarelli for helpful discussions on the Cramér-Rao type bound.
+Also, the author would like to thank Prof.
+Macro Tomamichel and Dr.
+Hu Yanglin for helpful discussions for the relation with full tomography and informing the references [27, 34].
 Appendix A Evaluation of mean square error: Proof of Theorem 2
@@
 $=$ $\mathrm{Tr}\rho^{\otimes n}\Big{(}\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{)}^{2}$
-$=$ $\frac{1}{n^{2}}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{n^{2}}(\log d_{n})^{2},$ (58)
+$=$ $\frac{1}{n^{2}}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{n^{2}}(\log d_{n,d})^{2},$ (60)
 where $(a)$ follows from [3, Lemma 8]. The combination of (58), (59), and (60) implies (22).
@@
 $=$ $\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)(\frac{1}{\sqrt{n}}\log P_{\rho,\sigma}^{(n)}(\bm{\lambda},j))^{2}$
-$=$ $\frac{1}{n}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\leq\frac{1}{n}(\log d_{n})^{2}$
+$=$ $\frac{1}{n}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\leq\frac{1}{n}(\log d_{n,d})^{2}$
 $\leq$ $\frac{1}{n}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}\to 0.$ (62)
@@
 We have the pinching inequality:
-$\rho^{\prime}\leq d_{n}\Gamma_{\sigma,n}(\rho^{\prime}).$ (63)
+$\rho^{\prime}\leq d_{n,d}\Gamma_{\sigma,n}(\rho^{\prime}).$ (65)
 In fact, applying the matrix monotone function to (65), for $\alpha\in(0,1)$, we have
-$d_{n}^{\alpha-1}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
-$=$ $\Big{(}d_{n}\Gamma_{\sigma,n}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}^{\alpha-1}$
+$d_{n,d}^{\alpha-1}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
+$=$ $\Big{(}d_{n,d}\Gamma_{\sigma,n}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}^{\alpha-1}$
 $\leq$ $\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}.$ (66)
@@
 $\cdot\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}$
-$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
+$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n,d}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
 $\cdot((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
-$=$ $d_{n}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
-$=$ $d_{n}^{1-\alpha}(\mathrm{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n},$ (65)
+$=$ $d_{n,d}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
+$=$ $d_{n,d}^{1-\alpha}(\mathrm{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n},$ (67)
 where Step $(a)$ follows from (66).
@@
 $=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}(\sigma^{\otimes n})^{\alpha}e^{n\alpha R}$
-$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
-$=$ $d_{n}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)}.$ (66)
+$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n,d}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
+$=$ $d_{n,d}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)}.$ (68)
 Here, $(b)$ follows from $\kappa_{\sigma,n}(\rho^{\otimes n})\leq I$, and $(c)$ follows from (67).
@@
 $+\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\rho^{\otimes n})\leq e^{-nr}\}$
-$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
-$=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
-$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\mathrm{Tr}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
-$=$ $(\mathrm{Tr}(\sigma^{-\frac{\alpha}{2(1+\alpha)}}\rho\sigma^{-\frac{\alpha}{2(1+\alpha)}})^{1+\alpha}e^{-\alpha(R-r)})^{n}+d_{n}e^{-nr}$
-$=$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n}e^{-nr}.$ (67)
+$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
+$=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
+$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\mathrm{Tr}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
+$=$ $(\mathrm{Tr}(\sigma^{-\frac{\alpha}{2(1+\alpha)}}\rho\sigma^{-\frac{\alpha}{2(1+\alpha)}})^{1+\alpha}e^{-\alpha(R-r)})^{n}+d_{n,d}e^{-nr}$
+$=$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n,d}e^{-nr}.$ (69)
 $(d)$ follows from the monotonicity of sandwitch Rényi relative entropy.

### v4

@@
 Our proposed estimator unifiedly works under both settings.
+[…]
 relative entropy; sample complexity; Cramer-Rao type bound; Schur composition
-I Introduction
+1 Introduction
 Distinguishing a given known state $\sigma$ from another state $\rho$ is a fundamental task in quantum information.
@@
 Therefore, this paper studies how to estimate the quantum relative entropy $D(\rho\|\sigma)$ with $n$ copies of the unknown state $\rho$ when $\sigma$ is known.
+In fact, when we intend to prepare a specific state $\sigma$, it is crucial to identify the difference between the known target state $\sigma$ and the unknown actual real state $\rho$.
+In this situation, we often focus on the fidelity $F(\rho,\sigma):=\mathrm{Tr}|\sqrt{\rho}\sqrt{\sigma}|$ for this aim.
+Many papers [14, 15, 16, 17, 18] studied this problem when $\sigma$ is a pure state.
+However, no paper focuses on this topic when $\sigma$ is a mixed state.
+Instead of the fidelity, one can evaluate the difference between $\sigma$ and $\rho$ by using the quantum relative entropy $D(\rho\|\sigma)$ because of the bounding inequality $-2\log F(\rho,\sigma)\leq D(\rho\|\sigma)$ [19, Eq.(3.49)].
+That is, our problem works for evaluating this difference.
 The estimation of $D(\rho\|\sigma)$ with the unknown state $\rho$ is formulated as a special case of the estimation of a function output with nuisance parameters.
 This general problem setting has been studied in various researches [20, 21, 22, 23, 24, 25].
-The reference [19, Section III] addressed the estimation problem of $D(\rho\|\sigma)$ as a special case, and showed that the estimation error is lower bounded by the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing [20, 21] while it did not show its attainability.
-Under a general setting, the reference [16, Section 4.3] showed that the optimal performance of the estimation precision is given by the Cramér-Rao type bound $C$ with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case.
-The meaning of the Cramér-Rao type bound $C$ is the following; when $n$ copies of the unknown state $\rho$ is prepared under a fixed model, the mean square of our best estimator behaves as $\frac{C}{n}+o(\frac{1}{n})$.
+The reference [25, Section III] addressed the estimation problem of $D(\rho\|\sigma)$ as a special case, and showed that the estimation error is lower bounded by the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing [26, 27] while it did not show its attainability.
+Under a general setting, the reference [22, Section 4.3] showed that the optimal precision of the estimation is given by the Cramér-Rao type bound $C$ with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case.
+The meaning of the Cramér-Rao type bound $C$ is the following: when $n$ copies of the unknown state $\rho$ is prepared under a fixed model, the mean square of our best estimator behaves as $\frac{C}{n}+o(\frac{1}{n})$.
 When we have no nuisance parameter, the recent paper [28] showed the existence of an estimator whose mean square error (MSE) equals $\frac{C}{n}$ without additional term under special models.
 However, such an estimator does not exist in general.
-In this paper, as the first step, using the results for estimation with nuisance parameters in [16, Section 4.3], we show that the Cramér-Rao type bound $C$ of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which was obtained as a lower bound in [19].
+In this paper, as the first step, using the results for estimation with nuisance parameters in [22, Section 4.3], we show that the Cramér-Rao type bound $C$ of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which was obtained as a lower bound in [25].
 In practice, it is necessary to find an estimator to achieve the MSE $\frac{C}{n}+o(\frac{1}{n})$.
@@
 However, this method requires an adaptive choice of measurement.
-One typical method for this approach is a two-step estimator, which is composed of the pre-estimation stage and the local optimal estimation.
+A typical method is the two-step estimator, consisting of a pre-estimation stage and a local optimal estimation stage.
 When we have $n$ copies of the unknown state $\rho$, theoretically, in order to achieve the optimal bound, it is sufficient to allocate $\sqrt{n}$ copies to the pre-estimation stage [29, 21].
@@
 In addition, the above adaptive approach does not work when the dimension $d$ of the quantum system increases in addition to the number $n$ of copies.
-In summary, considering the finite size effect of our estimation, we need to find an estimator of the quantum relative entropy $D(\rho\|\sigma)$ to satisfy the following conditions.
+In summary, to account for finite-size effects in our estimation, we need to find an estimator of the quantum relative entropy $D(\rho\|\sigma)$ to satisfy the following conditions.
 It asymptotically attains the Cramér-Rao type bound $V(\rho\|\sigma)$ in the sense of the order $\frac{1}{n}$ when the dimension $d$ is fixed.
-The difference between its mean square error and the Cramér-Rao type bound is evaluated with the number $n$ of copies.
-Further, the estimator works well even when the dimension $d$ of the quantum system increases in addition to the number $n$ of the prepared copies.
+The difference between its mean square error and the Cramér-Rao type bound is evaluated in terms of the number $n$ of copies.
+Furthermore, the estimator performs well even when the dimension $d$ of the quantum system increases in addition to the number $n$ of the prepared copies.
 As the second step of this paper, to satisfy the above requirements, we propose an alternative estimation method by using Schur transform [32].
-Then, we evaluate its mean square error and clarify its difference from the Cramér-Rao type bound with a finite number $n$.
+Then, we evaluate its mean square error and compare it to the Cramér-Rao type bound with a finite number $n$.
 In addition, we evaluate the tail probability.
 As the third step, similar to the papers [6, 7], we study the case when the dimension $d$ increases when $n$ increases.
-Unlike the estimation of von Neumann entropy, the range of the quantum relative entropy nor the relative varentropy is not bounded even with fixed dimension $d$.
+Unlike the estimation of von Neumann entropy, the range of the quantum relative entropy or the relative varentropy is not bounded even with fixed dimension $d$.
 Therefore, to study the sample complexity, we need to restrict the possible range of $\rho$ and $\sigma$.
@@
 Our result contains the case studied in the papers [6, 7].
-In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ for the Empirical Young Diagram (EYD) algorithm, which is given as the entropy of the diagonal state determined by the Young indecies.
+In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ for the Empirical Young Diagram (EYD) algorithm, which is given as the entropy of the diagonal state determined by the Young indices.
 Although this case improves the sample complexity over the simple application of the full tomography, the improvement is incremental.
-The paper [7] also showed that the lower bound of the sample complexity is $O(d)$ without this restriction.
+The paper [7] also showed a lower bound $O(d)$ for the sample complexity, without restricting our algorithm to EYD algorithm.
 In this special case, we show that our measurement can be restricted to Schur sampling.
 The remaining part of this paper is organized as follows.
-First, Section II derives the Cramér-Rao type bound for the estimation of quantum relative entropy.
-Section III presents our estimator for quantum relative entropy, and its performance.
+First, Section 2 derives the Cramér-Rao type bound for the estimation of quantum relative entropy.
+Section 3 presents our estimator for quantum relative entropy, and its performance.
 Then, this section shows that this estimator asymptotically attains the Cramér-Rao type bound when $d$ is fixed.
-Section IV studies the sample complexity of our method when $d$ increases.
-Section V studies the time complexity of our method when $d$ increases.
-Appendices are devoted to the proofs of the results presented in Section III.
-II Cramér-Rao type bound
-To clarify the minimum mean square error with the first order $\frac{1}{n}$, we discuss the Cramér-Rao type bound $C$ with nuisance parameters, which is mathematically formulated as a lower bound of MES under the local unbiasedness condition.
-The reference [19, (3.24), (3.52)] showed that the varentropy $V(\rho_{*}\|\sigma)$ is a lower bound under the estimation of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$.
+Section 4 studies the sample complexity of our method when $d$ increases.
+Section 5 studies the time complexity of our method when $d$ increases.
+Appendices are devoted to the proofs of the results presented in Section 3.
+2 Cramér-Rao type bound
+To clarify the minimum mean square error with the first order $\frac{1}{n}$, we discuss the Cramér-Rao type bound $C$ with nuisance parameters, which is mathematically formulated as a lower bound of MSE under the local unbiasedness condition.
+The reference [25, (3.24), (3.52)] showed that the varentropy $V(\rho_{*}\|\sigma)$ is a lower bound under the estimation of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$.
 That is, they showed that
@@
 For this aim, we review the Cramér-Rao type bound $C$ for a general function $f(\rho)$ in the full parameter model under a $d$-dimensional system as follows.
-We choose a parametrization $\rho_{\theta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{d^{2}-1})\in\mathbb{R}^{d^{2}-1}$ such that the first parameter gives the value of the function to be estimated and the other parameters are orthogonal to the first parameter at $\rho=\rho_{*}$.
+We choose a parameterization $\rho_{\theta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{d^{2}-1})\in\mathbb{R}^{d^{2}-1}$ such that the first parameter gives the value of the function to be estimated and the other parameters are orthogonal to the first parameter at $\rho=\rho_{*}$.
 That is, our parameter needs to satisfy the following conditions.
@@
 for $j=2,\ldots,d^{2}-1$. Then, we can calculate the Cramér-Rao type bound, i.e., the minimum mean square error under the local unbiasedness condition even under the full parameter model.
-Proposition 1 ([16, Theorem 5.3]).
+Proposition 1 ([22, Theorem 5.3]).
 Assume the conditions (2)–(6). The Cramér-Rao type bound for the estimation of $f(\rho_{\theta})$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}$ at $\theta=\theta_{0}$.
@@
 $\mathrm{Tr}X_{j}L^{1}$ $=0$ (9)
-for $j=2,\ldots,d^{2}-1$. Then, we choose our parametrization $\rho_{\theta}$ as
+for $j=2,\ldots,d^{2}-1$. Then, we choose our parameterization $\rho_{\theta}$ as
 $\rho_{\theta}:=\rho_{*}+(\theta^{1}-\theta^{1}_{0})X_{1}+\sum_{j=2}^{d^{2}-1}\theta^{j}X_{j}.$ (10)
@@
 $\frac{\partial D(\rho_{\theta}\|\sigma)}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
-$=$ $\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}\rho_{*}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
+$=$ $\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}(\log\rho_{*}-\log\sigma)$
+$+\mathrm{Tr}\rho_{*}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
 =$\stackrel{{\scriptstyle(a)}}{{=}}$ $\mathrm{Tr}X_{1}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}X_{1}$
 =$\stackrel{{\scriptstyle(b)}}{{=}}$ $\mathrm{Tr}X_{1}L^{1}+\mathrm{Tr}X_{1}D(\rho_{*}\|\sigma)I+\mathrm{Tr}X_{1}\stackrel{{\scriptstyle(c)}}{{=}}1.$ (11)
-Here, $(b)$ and $(c)$ follows from (7) and (8) and the relation $\langle L^{1},L^{1}\rangle_{\rho_{*}}=V(\rho_{*}\|\sigma)$, respectively. In addition, $(a)$ can be shown as follows.
-$\mathrm{Tr}\rho_{\theta_{0}}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}\stackrel{{\scriptstyle(a)}}{{=}}\mathrm{Tr}\int_{0}^{1}\rho_{\theta_{0}}^{s}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}\rho_{\theta_{0}}^{1-s}ds$
+Here, $(b)$ and $(c)$ follow from (7) and the pair of (8) and the relation $\langle L^{1},L^{1}\rangle_{\rho_{*}}=V(\rho_{*}\|\sigma)$, respectively. In addition, $(a)$ can be shown as follows.
+$\mathrm{Tr}\rho_{\theta_{0}}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
+=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\mathrm{Tr}\int_{0}^{1}\rho_{\theta_{0}}^{s}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}\rho_{\theta_{0}}^{1-s}ds$
 $=$ $\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}=\mathrm{Tr}X_{1},$ (12)
-where $(a)$ follows from [28, Eq. (6.32)].
+where $(a)$ follows from [19, Eq. (6.32)].
 Using (9) and a similar discussion, we have
-$\frac{\partial D(\rho_{\theta}\|\sigma)}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}=\mathrm{Tr}X_{j}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}X_{j}=0.$ (13)
+$\frac{\partial D(\rho_{\theta}\|\sigma)}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}=$ $\mathrm{Tr}X_{j}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}X_{j}$
+$=$ $0$ (13)
 Then, (12) and (13) imply the conditions (3) and (4).
@@
 which implies the condition (6).
-Therefore, since all the conditions (2) – (6) hold, due to Proposition 1, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}=V(\rho_{*}\|\sigma)^{-1}$ at $\theta=\theta_{0}$.
+Therefore, since all the conditions (2) – (6) hold, due to Proposition 1, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}=V(\rho_{*}\|\sigma)$ at $\theta=\theta_{0}$.
 That is, the minimum mean square error under the local unbiasedness condition is $V(\rho_{*}\|\sigma)$.
@@
 Thus, the quality in (1) holds.
-III Estimator based on Schur transform
+3 Estimator based on Schur transform
 Similar to the papers [34, 3, 8, 9, 10, 11, 12, 13, 5, 35, 36], we employ Schur duality of ${\cal H}^{\otimes n}$.
@@
 Although many references [36, 37, 38] define Young index as monotone-decreasing non-negative integers, we define it in the opposite way for notational convenience.
-We denote the set of Young indices $\bm{\lambda}$ with the condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [31, Section 6.2], we have
+We denote the set of Young indices $\bm{\lambda}$ with the condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [36, Section 6.2], we have
 ${\cal H}^{\otimes n}=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}\otimes{\cal V}_{\bm{\lambda}}.$ (15)
@@
 Then, for states $\sigma$, $\rho$, we have
-$\rho^{\otimes n}$ $=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (16)
-$\sigma^{\otimes n}$ $=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (17)
+$\rho^{\otimes n}$ $=\sum_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (16)
+$\sigma^{\otimes n}$ $=\sum_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (17)
 where $\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits}$ is the completely mixed state on ${\cal V}_{\bm{\lambda}}$.
+In the above equations, we consider that matrices on the subspace ${\cal U}_{\bm{\lambda}}\otimes{\cal V}_{\bm{\lambda}}$ are matrices on the whole space ${\cal H}^{\otimes n}$.
+In the first step, we apply the Schur transform $U_{Schur}$, which realizes a basis change to block-diagonalize $U^{\otimes n}$ for an arbitrary unitary $U$.
+It is known that the Schur transform $U_{Schur}$ can be efficiently done [32].
 We choose a basis $\{|u_{\bm{\lambda},j}\rangle\}_{j=1}^{d_{\bm{\lambda}}}$ of ${\cal U}_{\bm{\lambda}}$ such that $|u_{\bm{\lambda},j}\rangle$ is an eigenvector of $\sigma_{\bm{\lambda}}$.
@@
 To estimate the quantum relative entropy $D(\rho\|\sigma)$, we employ the estimate $x_{\bm{\lambda},j}$ defined as $-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$.
-This measurement can be done by Schur transform efficiently [26].
+Our estimation protocol is summarized as Protocol 1.
+Protocol 1 Estimation of relative entropy
+  1. Step 1.
+    We apply the Schur transform $U_{Schur}$.
+  2. Step 2.
+    We measure the system by using the measurement $M=\{M_{\bm{\lambda},j}\}_{\bm{\lambda},j}$.
+  3. Step 3.
+    We set our estimate to be $x_{\bm{\lambda},j}:=-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$.
 To evaluate the mean square error of this estimator, we denote the dimension of $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$ by $d_{n,d}$.
-Since as shown in [31, (6.16) and (6.18)], we have
+Since as shown in [36, (6.16) and (6.18)], we have
 $\dim{\cal U}_{\bm{\lambda}}\leq(n+1)^{d(d-1)/2},\quad|Y_{d}^{n}|\leq(n+1)^{d-1},$ (20)
@@
 $d_{n,d}\leq(n+1)^{(d+2)(d-1)/2}$ (21)
-holds. The mean square error of this estimator is evaluated as follows.
+holds.
+The reason of our choices of the POVM $M_{\bm{\lambda},j}$ and the estimate $x_{\bm{\lambda},j}$ is the following. The reference [34] essentially showed the relation
+$D(\rho^{\otimes n}\|\sigma^{\otimes n})-\frac{d(d-1)}{2}\log(n+1)$
+$\leq$ $D_{M}(\rho^{\otimes n}\|\sigma^{\otimes n})\leq D(\rho^{\otimes n}\|\sigma^{\otimes n}),$ (22)
+where
+$D_{M}(\rho^{\otimes n}\|\sigma^{\otimes n})$
+::=:= $\sum_{\bm{\lambda},j}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}(\log\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}-\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}).$
+This relation shows that the POVM $M_{\bm{\lambda},j}$ almost preserves the information with respect to the relative entropy $D(\rho^{\otimes n}\|\sigma^{\otimes n})$.
+Further, since the Shannon entropy $H(P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)):=-\sum_{\bm{\lambda},j}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}\log\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}$ satisfies
+$0\leq H(P_{\rho,\sigma}^{(n)}(\bm{\lambda},j))\leq\log d_{n,d},$ (23)
+the the combination of (21) and (22) implies
+$D(\rho^{\otimes n}\|\sigma^{\otimes n})-(d+1)(d-1)\log(n+1)$
+$\leq$ $-\sum_{\bm{\lambda},j}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$
+$\leq$ $D(\rho^{\otimes n}\|\sigma^{\otimes n}).$ (24)
+That is, the average of $-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$ is close to the value $D(\rho\|\sigma)$ to be estimated.
+This is the reason why we choose the estimate $x_{\bm{\lambda},j}$.
+The mean square error of this estimator is evaluated as follows.
 Theorem 2.
@@
 This theorem is proved in Appendix A.
-The difference between the upper bound (22) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}\log d_{n,d}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n,d}$.
+The difference between the upper bound (25) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}(\log d_{n,d})^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n,d}$.
 Using (21), this different is upper bounded by $\frac{1}{n^{2}}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\frac{(d+2)(d-1)}{2}\log(n+1)$.
@@
 For any states $\rho$, $\sigma$, and $R,r>0$, we have
-$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $d_{n,d}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (27)
-$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n,d}e^{-nr}.$ (28)
+$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $(d_{n,d})^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (30)
+$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}$
+$+d_{n,d}e^{-nr}.$ (31)
 In (30), $\alpha\in(0,1)$, and in (31), $\alpha>0$.
 Theorem 4 guarantees that $\delta_{\rho,\sigma}^{(n),+,\epsilon}$ and $\delta_{\rho,\sigma}^{(n),-,\epsilon}$ go to zero exponentially. This property is called the large deviation.
-IV Analysis with increasing $d$
+4 Analysis with increasing $d$
 Similar to [6, 7], we consider the case when the dimension $d$ increases.
@@
 In this case, due to Theorem 2, the Cramér-Rao type bound can be attained because the second term of RHS in (25) goes to zero.
-Now, we consider when the error goes to zero under the condition $d\to\infty$.
+We now discuss the behavior of $n$ with respect to $d$ that allows the error to converge to zero as $d\to\infty$, rather than assuming $n=O(d^{4+\delta})$.
 We denote the $d$-dimensional system by ${\cal H}_{d}$.
@@
 $=$ $\|\log\rho-(\mathrm{Tr}\rho\log\rho)-\log\sigma_{d}+(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}$
-$\leq$ $\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}+\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}.$ (30)
+$\leq$ $\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}$
+$+\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}.$ (33)
 Using the property of the norm $\|~{}\|_{\rho}$, we have
-$\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}^{2}=\mathrm{Tr}\rho(\log\rho-(\mathrm{Tr}\rho\log\rho))^{2}$
-$=$ $(\mathrm{Tr}\rho(\log\rho)^{2})-(\mathrm{Tr}\rho\log\rho)^{2}\leq(\mathrm{Tr}\rho(\log\rho)^{2})\stackrel{{\scriptstyle(a)}}{{\leq}}(\log d)^{2},$ (31)
+$\|\log\rho-(\mathrm{Tr}\rho\log\rho)\|_{\rho}^{2}$
+$=$ $\mathrm{Tr}\rho(\log\rho-(\mathrm{Tr}\rho\log\rho))^{2}$
+$=$ $(\mathrm{Tr}\rho(\log\rho)^{2})-(\mathrm{Tr}\rho\log\rho)^{2}$
+$\leq$ $(\mathrm{Tr}\rho(\log\rho)^{2})\stackrel{{\scriptstyle(a)}}{{\leq}}(\log d)^{2},$ (34)
 where $(a)$ follows from [3, Lemma 8]. Similarly, we have
-$\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}^{2}=\mathrm{Tr}\rho(\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d}))^{2}$
-$=$ $(\mathrm{Tr}\rho(\log\sigma_{d})^{2})-(\mathrm{Tr}\rho\log\sigma_{d})^{2}\leq(\mathrm{Tr}\rho(\log\sigma_{d})^{2})\leq(td)^{2}.$ (32)
+$\|\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d})\|_{\rho}^{2}$
+$=$ $\mathrm{Tr}\rho(\log\sigma_{d}-(\mathrm{Tr}\rho\log\sigma_{d}))^{2}$
+$=$ $(\mathrm{Tr}\rho(\log\sigma_{d})^{2})-(\mathrm{Tr}\rho\log\sigma_{d})^{2}$
+$\leq$ $(\mathrm{Tr}\rho(\log\sigma_{d})^{2})\leq(td)^{2}.$ (35)
 Combining (33), (34), and (35), we obtain the condition (32). ∎
@@
 When the condition (32) holds and $n=cd^{2}$, we have
-$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon}+\delta_{\rho,\sigma}^{(n),-,\epsilon}\leq$ $\frac{1}{\epsilon^{2}}\Big{(}\frac{\sqrt{c_{0}}}{\sqrt{c}}+\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}\Big{)}^{2}$
+$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon}+\delta_{\rho,\sigma}^{(n),-,\epsilon}$
+$\leq$ $\frac{1}{\epsilon^{2}}\Big{(}\frac{\sqrt{c_{0}}}{\sqrt{c}}+\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}\Big{)}^{2}$
 $\leq$ $\frac{(\sqrt{c_{0}}+4)^{2}}{c\epsilon^{2}}.$ (36)
@@
 Now, we compare the sample complexity of our method with the sample complexity of the simple application of full tomography [33, 39].
-In particular, we employ the error analysis on the trance norm [27] rather than other error criteria because the trance norm is most useful to evaluate the error of the relative entropy.
+In particular, we employ the error analysis on the trace norm [33] rather than other error criteria because the trace norm is most useful to evaluate the error of the relative entropy.
 Their method achieves the estimation error
@@
 That is, our extension to the estimation of the relative entropy realizes significant improvement in the sample complexity over the above method.
+In fact, the paper [39] also considers the sample complexity when Bures $\chi^{2}$-divergence and Bures distance are adopted.
+However, the evaluation of the error under these criteria does not fit well with the above method.
+Moreover, it is challenging to find a good upper bound of the estimation errors of $S(\hat{\rho})-S(\rho)$ and $D(\hat{\rho}\|\sigma)-D(\rho\|\sigma)$ by using the estimation error based on these criteria.
 Proof of Theorem 6.
-We choose $s\in(0,1)$. Using Weyl’s dimension formula [32, Theorem 7.1.7], we have
+We choose $s\in(0,1)$. Using Weyl’s dimension formula [38, Theorem 7.1.7], we have
 $\log\dim{\cal U}_{\bm{\lambda}}=\log\prod_{1\leq i<j\leq d}\frac{j-i+\lambda_{j}-\lambda_{i}}{j-i}$
-$=$ $\!\!\sum_{1\leq i<j\leq d}\!\!\log\Big{(}1+\frac{\lambda_{j}-\lambda_{i}}{j-i}\Big{)}=\sum_{l=1}^{d}\sum_{i=1}^{d-l}\log\Big{(}1+\frac{\lambda_{l+i}-\lambda_{i}}{l}\Big{)}$
+$=$ $\!\!\sum_{1\leq i<j\leq d}\!\!\log\Big{(}1+\frac{\lambda_{j}-\lambda_{i}}{j-i}\Big{)}$
+$=$ $\sum_{l=1}^{d}\sum_{i=1}^{d-l}\log\Big{(}1+\frac{\lambda_{l+i}-\lambda_{i}}{l}\Big{)}$
 $=$ $\sum_{l=1}^{d}(d-l)\sum_{i=1}^{d-l}\frac{1}{d-l}\log\Big{(}1+\frac{\lambda_{l+i}-\lambda_{i}}{l}\Big{)}$
@@
 The idea of this derivation comes from the proof of (12) given in Appendix IV of the paper [40], which gives the security analysis in information-theoretic security.
-V Analysis on time complexity
+5 Analysis on time complexity
 To implement our estimator, we need to apply suitable unitary gates before the measurement based on the computation basis.
@@
 For a general density $\sigma$, we need to apply a unitary diagonalizing the density $\sigma$ to each system ${\cal H}$ before the following procedure.
-First, we apply the Schur transform $U_{Schur}$, which realizes a basis change to block-diagonalize $U^{\otimes n}$ for an arbitrary unitary $U$.
-The references [26, 36, 37, 38] proposed its quantum algorithm.
-We evaluate its prevision as follows.
+The references [32, 41, 42, 43] proposed quantum algorithms for the Schur transform $U_{Schur}$.
+We evaluate its precision as follows.
 Given a required error $\epsilon>0$ and a target unitary $V$, we say that a unitary $U$ is an $\epsilon$-approximate of $V$ when any pure state $|\psi\rangle$ satisfies
@@
 $e(\bm{\lambda})$ $:=\prod_{j>i}\frac{\lambda_{j}-\lambda_{i}-i+j}{\lambda_{j}+j-i}<1,$ (53)
-where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [33, Eq. (2.72)]. Using [28, Theorem 2.5], we have
+where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [37, Eq. (2.72)]. Using [19, Theorem 2.5], we have
 $e^{nH(\frac{\bm{\lambda}}{n})}(n+1)^{-(d-1)}\leq\frac{n!}{\bm{\lambda}!}\leq e^{nH(\frac{\bm{\lambda}}{n})}.$ (54)
@@
 $-\log e(\bm{\lambda})=\sum_{j>i}\log\frac{\lambda_{j}+j-i}{\lambda_{j}-\lambda_{i}-i+j}$
-$=$ $\sum_{j>i}\log(1+\frac{\lambda_{i}}{\lambda_{j}-\lambda_{i}+j-i})\leq\sum_{j>i}\log(1+\frac{\lambda_{i}}{j-i})$
+$=$ $\sum_{j>i}\log(1+\frac{\lambda_{i}}{\lambda_{j}-\lambda_{i}+j-i})$
+$\leq$ $\sum_{j>i}\log(1+\frac{\lambda_{i}}{j-i})$
 $=$ $\sum_{l=1}^{d}(d-l)\sum_{i=1}^{d-l}\frac{1}{d-l}\log\Big{(}1+\frac{\lambda_{i}}{l}\Big{)}$
 $\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\sum_{i=1}^{d-l}\frac{1}{d-l}\frac{\lambda_{i}}{l}\Big{)}$
-$\leq$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\frac{n}{(d-l)l}\Big{)}\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}},$ (52)
+$\leq$ $\sum_{l=1}^{d}(d-l)\log\Big{(}1+\frac{n}{(d-l)l}\Big{)}$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}},$ (55)
 where $(a)$ follows from Jensen inequality for the concave function $x\mapsto\log(1+x)$, and $(b)$ follows from (44) in the same way as (43).
@@
 When $n=cd^{2}$, $\lim_{n\to\infty}\frac{1}{n}(d\log(n+1)-\sum_{l=1}^{d}\frac{(d-l)^{1-s}n^{s}}{sl^{s}})\leq\frac{4}{\sqrt{c}}$.
-Given an arbitrary number $\epsilon>0$, we choose $c=\frac{16}{\epsilon}$. When $n=cd^{2}$, the limit of the upper bound in (53) can be bounded by $\epsilon$. Then, we define
+Given an arbitrary number $\epsilon>0$, we choose $c=\frac{16}{\epsilon^{2}}$. When $n=cd^{2}$, the limit of the upper bound in (56) can be bounded by $\epsilon$. Then, we define
 $\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}$ $:=P_{\rho,\sigma}^{(n),\epsilon_{1}}\{(\bm{\lambda},j):x_{\bm{\lambda},j}^{*}-D(\rho\|\sigma)>\epsilon\}$ (57)
@@
 Thus, when $n=cd^{2}$, using (37), we have
-$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}+\delta_{\rho,\sigma}^{(n),-,\epsilon,\epsilon_{1}}\leq\frac{(\sqrt{c_{0}}+4)^{2}}{c(\epsilon-\frac{4}{\sqrt{c}})^{2}}+\epsilon_{1}^{2}.$ (56)
+$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\delta_{\rho,\sigma}^{(n),+,\epsilon,\epsilon_{1}}+\delta_{\rho,\sigma}^{(n),-,\epsilon,\epsilon_{1}}$
+$\leq$ $\frac{(\sqrt{c_{0}}+4)^{2}}{c(\epsilon-\frac{4}{\sqrt{c}})^{2}}+\epsilon_{1}^{2}.$ (59)
 Hence, when we choose $\epsilon_{1}=\sqrt{\epsilon^{\prime}/2}$ and $c=\min(\frac{64}{\epsilon^{2}},\frac{8(4+\sqrt{c_{0}})^{2}}{\epsilon^{2}\epsilon^{\prime}})$ for any $\epsilon,\epsilon^{\prime}>0$, we have
@@
 The paper [7] also mentions that there exists an algorithm for Schur transform $U_{Schur}$ whose time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$ when $n=O(d^{2})$ [44].
-First, they employ the algorithm for weak Schur sampling given in [40, Section 4.2.2].
+First, they employ the algorithm for weak Schur sampling given in [45, Section 4.2.2].
 Although this method employs the quantum Fourier transform the permutation group $\mathfrak{S}_{n}$, they proposed to use the algorithm by [46] for this part.
@@
 When $n=d^{2}$, it equals $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$.
-This implementation of weak Schur sampling was mentioned earlier in [42, Section 1.1.2].
-VI Discussion
+This implementation of weak Schur sampling was mentioned earlier in [47, Section 1.1.2].
+Remark 9.
+Recently, the paper [48] studied the estimation of relative entropy with two unknown densities $\rho$ and $\sigma$ when an estimator of the state is given and the estimator converges to the true density.
+The analysis [48] is different from ours in the following points.
+(i) We assume that $\sigma$ is known, but the paper [48] assumes that it is unknown.
+(ii) We discuss the optimality of our estimator in various settings.
+Under the case with fixed $d$, we optimize the MSE.
+When $d$ increases, we discuss the sample complexity.
+But, when an estimator of the state is given, the paper [48] discusses the behavior of the estimator of relative entropy by using the behavior of the estimator.
+In addition, the paper [48] mentioned that the limiting distribution of the normalized estimator is the normal distribution.
+When $d$ is fixed, the discussion explains how this problem is reduced to a special case of the parameter estimation with nuisance parameters.
+The paper [21] also showed that the limiting distribution of the normalized estimator is the normal distribution in the parameter estimation with nuisance parameters.
+6 Discussion
 We have studied the estimation of quantum relative entropy $D(\rho\|\sigma)$ when the second argument $\sigma$ is known.
@@
 The author is supported in part by the National Natural Science Foundation of China (Grant No.
-62171212).
+62171212) and the General R&D Projects of 1+1+1 CUHK-CUHK(SZ)-GDST Joint Collaboration Fund (Grant No.
+GRDP2025-022).
 The author is grateful for Professor Kun Fang to a helpful discussion.
@@
 We choose an orthogonal basis $\{|v_{\bm{\lambda},j^{\prime}}\}_{j^{\prime}}$ of ${\cal V}_{\bm{\lambda}}$.
+We use the methods [3, 26, 49]. Then, we have the following calculation.
+$x_{\bm{\lambda},j}-D(\rho\|\sigma)=-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}-D(\rho\|\sigma)$
+$=$ $-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log d_{\bm{\lambda}}^{\prime})$
+$-D(\rho\|\sigma)$
+$=$ $-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
+$-\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma)$
+$-\frac{1}{n}(\log d_{\bm{\lambda}}^{\prime}+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)$
+$=$ $\frac{1}{n}(-\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
+$+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma)$
+$-\frac{1}{n}\log\langle u_{\bm{\lambda},j}|\rho_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle.$ (61)
+Therefore, we have
 $\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)(x_{\bm{\lambda},j}-D(\rho\|\sigma))^{2}$
-$=$ $\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)\Big{(}\frac{1}{n}(-\log\langle u_{\bm{\lambda},j}|\sigma_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle-D(\rho\|\sigma)\Big{)}^{2}$
+$=$ $\sum_{(\bm{\lambda},j)}P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)\Big{(}\frac{1}{n}(-\log\langle u_{\bm{\lambda},j}|\sigma_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle$
+$-D(\rho\|\sigma)\Big{)}^{2}$
 $=$ $\sum_{(\bm{\lambda},j,j^{\prime})}\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
@@
 and
-$\Big{\|}-\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{\|}_{\rho^{\otimes n}}$
+$\Big{\|}-\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{\|}_{\rho^{\otimes n}}^{2}$
 $=$ $\mathrm{Tr}\rho^{\otimes n}\Big{(}\frac{1}{n}\sum_{\bm{\lambda}}\log(\rho_{\bm{\lambda}}\otimes I_{{\cal V}_{\bm{\lambda}}})\Big{)}^{2}$
@@
 Appendix B Local asymptotic normality: Proof of Theorem 3
-We use the methods [3, 20, 44].
-$x_{\bm{\lambda},j}-D(\rho\|\sigma)=-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}-D(\rho\|\sigma)$
-$=$ $-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log d_{\bm{\lambda}}^{\prime})-D(\rho\|\sigma)$
-$=$ $-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
-$-\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma)$
-$-\frac{1}{n}(\log d_{\bm{\lambda}}^{\prime}+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)$
-$=$ $\frac{1}{n}(-\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$
-$+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma)$
-$-\frac{1}{n}\log\langle u_{\bm{\lambda},j}|\rho_{\bm{\lambda}}|u_{\bm{\lambda},j}\rangle.$ (61)
 The random variable $\sqrt{n}(-\frac{1}{n}(\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\sigma^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle+\log\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle)-D(\rho\|\sigma))$ has additional variable $j^{\prime}$.
 The joint distribution of $(\bm{\lambda},j,j^{\prime})$ is given as $\langle u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}|\rho^{\otimes n}|u_{\bm{\lambda},j},v_{\bm{\lambda},j^{\prime}}\rangle$.
-This random variable coincides with $\sqrt{n}(\frac{1}{n}(\log{\cal E}_{\sigma^{\otimes n}}(\rho^{\otimes n})-\log\sigma^{\otimes n})-D(\rho\|\sigma))$ under the state ${\cal E}_{\sigma^{\otimes n}}(\rho^{\otimes n})$ studied in Section V of [20].
+We define the pinching map $\Gamma_{\sigma,n}$ as
+$\Gamma_{\sigma,n}(\rho^{\prime})$
+::=:= $\sum_{\bm{\lambda}\in Y_{d}^{n},j}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))\rho^{\prime}$
+$\cdot(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))$ (65)
+for a general state $\rho^{\prime}$ on ${\cal H}^{\otimes n}$, where $P({\cal V}_{\bm{\lambda}})$ is the projection to ${\cal V}_{\bm{\lambda}}$.
+Then, this random variable coincides with $\sqrt{n}(\frac{1}{n}(\log\Gamma_{\sigma,n}(\rho^{\otimes n})-\log\sigma^{\otimes n})-D(\rho\|\sigma))$ under the state $\Gamma_{\sigma,n}(\rho^{\otimes n})$ studied in Section V of [26].
 The discussion in Section V of [26] showed that it converges to Gaussian distribution with variance $V(\rho\|\sigma)$ in probability as $n\to\infty$.
@@
 C.1 Preparation
-We define the pinching map $\Gamma_{\rho,n}$ as
-$\Gamma_{\sigma,n}(\rho^{\prime})$
-::=:= $\sum_{\bm{\lambda}\in Y_{d}^{n},j}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))\rho^{\prime}(|u_{\bm{\lambda},j}\rangle\langle u_{\bm{\lambda},j}|\otimes P({\cal V}_{\bm{\lambda}}))$ (63)
-for a general state $\rho^{\prime}$ on ${\cal H}^{\otimes n}$.
 Since the system ${\cal V}_{\bm{\lambda}}$ has no information, it is sufficient to handle the two states $\oplus_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}$ and $\oplus_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}$ over $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$.
@@
 Thus, we have
-$\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{\alpha}(\sigma^{\otimes n})^{1-\alpha}$
-$=$ $\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
-$=$ $\mathrm{Tr}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha}$
+$\mathrm{Tr}\big{(}\Gamma_{\sigma,n}(\rho^{\otimes n})\big{)}^{\alpha}(\sigma^{\otimes n})^{1-\alpha}$
+$=$ $\mathrm{Tr}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}$
+$=$ $\mathrm{Tr}\Big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha}$
 $=$ $\mathrm{Tr}\Big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}$
-$\cdot\big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}\big{)}$
-$=$ $\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
-$\cdot\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}$
-$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n,d}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
-$\cdot((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
-$=$ $d_{n,d}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
+$\cdot\Big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha-1}$
+$=$ $\mathrm{Tr}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}$
+$\cdot\Big{(}\Gamma_{\sigma,n}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha-1}$
+$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n,d}^{1-\alpha}\mathrm{Tr}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}$
+$\cdot\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$
+$=$ $d_{n,d}^{1-\alpha}\mathrm{Tr}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha}$
 $=$ $d_{n,d}^{1-\alpha}(\mathrm{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n},$ (70)
@@
 C.2 Proof of Theorem 4
-For $\alpha\in(0,1)$, we have
+Here, we define the projection $\{X>Y\}$ as the projection to the subspace spanned by the eigenvectors of $X-Y$ with positive eigenvalues. For $\alpha\in(0,1)$, we have
 $\sum_{(\bm{\lambda},j):x_{\bm{\lambda},j}<R}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}$
@@
 $\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\sigma^{\otimes n})>e^{-nR}\kappa_{\sigma,n}(\rho^{\otimes n})\}$
-$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{\alpha}e^{n\alpha R}$
+$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{\alpha}e^{n\alpha R}$
 $=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}(\sigma^{\otimes n})^{\alpha}e^{n\alpha R}$
-$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n,d}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
+$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $d_{n,d}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
 $=$ $d_{n,d}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)}.$ (71)
-Here, $(b)$ follows from $\kappa_{\sigma,n}(\rho^{\otimes n})\leq I$, and $(c)$ follows from (67).
+Here, $(b)$ follows from $\kappa_{\sigma,n}(\rho^{\otimes n})\leq I$, and $(d)$ follows from (70). The relation $(c)$ follows from
+$\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\sigma^{\otimes n})>e^{-nR}\kappa_{\sigma,n}(\rho^{\otimes n})\}$
+$=$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{e^{\alpha nR}(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}>(\kappa_{\sigma,n}(\rho^{\otimes n}))^{\alpha}\}$
+$\stackrel{{\scriptstyle(e)}}{{\leq}}$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}e^{n\alpha R}$
+$\cdot\{e^{\alpha nR}(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}>(\kappa_{\sigma,n}(\rho^{\otimes n}))^{\alpha}\}$
+[…]
+where $(e)$ follows from the following relation.
+$\mathrm{Tr}$ $\Big{[}(\kappa_{\sigma,n}(\rho^{\otimes n}))^{1-\alpha}$
+$\cdot\Big{(}(\kappa_{\sigma,n}(\rho^{\otimes n}))^{\alpha}-(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}e^{n\alpha R}\Big{)}$
+$\cdot\{e^{\alpha nR}(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}>(\kappa_{\sigma,n}(\rho^{\otimes n}))^{\alpha}\}\Big{]}\leq 0.$ (73)
 For $\alpha>0$, we have
@@
 $=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
-$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\mathrm{Tr}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
+$\stackrel{{\scriptstyle(f)}}{{\leq}}$ $\mathrm{Tr}((\sigma^{\otimes n})^{-\frac{\alpha}{2(1+\alpha)}\rho^{\otimes n}(\sigma^{\otimes n})^{-\frac{\alpha}{2(1+\alpha)}})^{1+\alpha}e^{-\alpha(R-r)n}}$
+$+d_{n,d}e^{-nr}$
 $=$ $(\mathrm{Tr}(\sigma^{-\frac{\alpha}{2(1+\alpha)}}\rho\sigma^{-\frac{\alpha}{2(1+\alpha)}})^{1+\alpha}e^{-\alpha(R-r)})^{n}+d_{n,d}e^{-nr}$
 $=$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n,d}e^{-nr}.\,$ (74)
-$(d)$ follows from the monotonicity of sandwitch Rényi relative entropy.
+where $(f)$ follows from the monotonicity of sandwitch Rényi relative entropy.